### v1

# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

Abstract
Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale results in high computational and memory costs. Recent studies have revealed that LLMs exhibit sparsity, providing the potential to reduce model size through pruning techniques. However, existing pruning methods typically follow a prune-then-finetune paradigm. Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, imposing a substantial computational burden to recover performance during finetuning. In this paper, we propose a novel paradigm that first applies regularization, then prunes, and finally finetunes. Based on this paradigm, we introduce DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs. By leveraging a small amount of data to regularize the components to be pruned, DReSS explicitly transfers the important information to the remaining parts of the model in advance. Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities. Experimental results demonstrate that DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
Machine Learning, ICML
1 Introduction
Large language models (LLMs) have achieved significant advancements across a wide range of tasks and domains, demonstrating their robust capabilities (Zhang et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Wu et al., 2024). However as the model size grows, the increased number of parameters results in substantial computational and memory demands. These factors severely hinder the deployment of LLMs in practical applications. Therefore, it is urgent to develop methods that can reduce model size while maintaining performance.
To address these challenges, several methods have been proposed, including pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Ma et al., 2023; An et al., 2024), quantization (Frantar et al., 2022; Xiao et al., 2023), knowledge distillation (Shridhar et al., 2022; Hsieh et al., 2023), and low-rank decomposition (Saha et al., 2023). In this work, we mainly focus on pruning, an efficient and highly generalizable method that can be combined with other model compression strategies. Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and structured pruning (Ma et al., 2023; An et al., 2024). Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) on the pruned model for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024). Moreover, the model obtained through structured pruning typically achieves much faster inference speed, while ensuring the same perplexity and accuracy as the model obtained through unstructured pruning.
Despite these advancements, existing structured pruning methods still have some limitations. They all follow the paradigm of first selecting channels or layers for pruning based on a designed metric, and then performing RFT (Chavan et al., 2024). However, they neglect the fact that important information can also exist in the pruned channels or layers (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance. Thus, in order to achieve satisfactory performance with the pruned model, a large amount of data must be introduced for RFT (Ma et al., 2023), which increases computational overhead. Moreover, they often lead to a collapse in model performance at high pruning ratios, limiting their effectiveness in reducing lantency and enhancing throughput.
To tackle these limitations, in this paper, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then removes those parameters, and finally performs RFT. To the best of our knowledge, we are the first to propose this paradigm. The regularization process explicitly transfers important information from the parts to be pruned to the remaining parts, thereby reducing the information loss caused by direct parameter removal. Based on this paradigm, we introduce DReSS, a Data-driven Regularized Structured Streamlining method for LLMs. The comparison of the proposed method with existing channel-wise pruning approaches is shown in Figure 1. Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, then (4) do RFT. First, we randomly selecting a small subset of data from widely used benchmark datasets, as the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced. Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization to these channels using the selected data. This shifts important information to the unpruned parts, significantly mitigating performance degradation caused by direct pruning and ensuring the model maintains strong performance even at high pruning ratios, thus enabling substantial inference speedup. Third, we prune the parameters that have been regularized. Finally, we do RFT on the pruned model using a small subset of samples selected in the first step. Extensive experiments demonstrate that, DReSS significantly surpasses existing pruning methods in terms of perplexity and accuracy across various pruning ratios, it also substantially improves inference speed. The main contributions of our work are summarized as follows:
  • Propose A New Paradigm: By first applying regularization, then pruning, and finally performing RFT, DReSS reduces information loss from direct parameter removal, enhancing the model’s language modeling ability and lowering the computational cost of RFT.
  • High Performance: DReSS outperforms existing structured pruning methods on both generation and zero-shot tasks. The pruned model shows a significant increase in throughput and a notable reduction in latency compared to the dense model.
  • Low Overhead: By using only a small amount of data during the regularization and optional RFT process, the pruning overhead is minimal.
Figure 1: A comparison between existing channel-wise structured pruning methods and DReSS. The deeper blue square represents greater performance impact; the red exclamation mark indicates the important information. The taller orange-yellow cylinder represents larger data volume. Here, L($\cdot$) represents the regularization, and S represents pruning. Above: Existing methods first select channels based on importance, followed by pruning and RFT on the pruned model. Below: DReSS first regularizes channels to minimize performance impact, prunes the channels, then do RFT with a little amount of data.
2 Related Works
2.1 Pruning Methods
The elimination of redundant weights, referred to as pruning, has been an effective approach for reducing the complexity of deep neural networks over the past several decades (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015; Wen et al., 2016; Louizos et al., 2017). Pruning methods can be broadly categorized into two types: unstructured pruning (Kurtic et al., 2022; Zhang et al., 2024; Xu et al., 2024) and structured pruning (Xia et al., 2023b; Gao et al., 2024b; Yang et al., 2024).
Unstructured pruning methods remove individual weights based on their importance. For instance, Magnitude pruning (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix. The main advantage is flexibility, enabling the removal of unimportant weights while retaining critical ones, allowing high sparsity in LLMs. However, its limitations include the need for specialized hardware for accelerated inference (Xia et al., 2023a), high sparsity for significant acceleration (Wang, 2020), and inability to perform RFT on downstream tasks.
Structured pruning methods, including channel-wise pruning (An et al., 2024) and layer-wise pruning (Men et al., 2024), can achieve significant inference speedup with a low sparsity. Channel-wise pruning methods typically design a metric to evaluate the importance of channels in the parameter matrix of LLMs and then prune the less important channels. For instance, SliceGPT (Ashkboos et al., 2024) adopts Principal Component Analysis (PCA), using the eigenvalues of the covariance matrix of activation values as a metric to preserve channels corresponding to larger eigenvalues. LLM Surgeon (van der Ouderaa et al., 2023) periodically updates model weights and structures, pruning more aggressively in the initial layers and less so in the intermediate layers. Layer-wise pruning methods, such as SLEB (Song et al., 2024), iteratively prune entire transformer layers by assessing their importance. However, these methods have a key limitation: even less important channels or layers may contain valuable information, and pruning them directly often leads to significant performance degradation. To address this, we investigate the possibility of ‘shifting’ important information from the pruned parts to the retained parts, which could substantially improve model performance.
2.2 Regularization
Regularization is widely employed in machine learning (Hoerl & Kennard, 1970; Poggio et al., 1987; Balestriero et al., 2022), such as feature selection (Tibshirani, 1996) and preventing model overfitting (Santos & Papa, 2022). The $\ell_{1}$-norm typically induces sparsity by driving certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd & Vandenberghe, 2004). Both of them can significantly alter the distribution pattern of the data (Han et al., 2015; Tao et al., 2023). Motivated by this observation, we propose a data-driven regularization approach for parameters designated for pruning. By applying a small amount of data, the regularization probably transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
3 Methodology
In this section, we provide a detailed introduction to DRess. The overall DReSS algorithm is summarized in Algorithm 1, our method consists of four main steps:
  • Data Selection: Select a small amount of data for pre-pruning regularization and post-pruning RFT.
  • Regularization: Applying regularization to the selected channels of weight matrices during both forward and backward propagations.
  • Pruing: Prune the regularized portions.
  • RFT: Perform RFT on the pruned model using the subset of data selected in the first step.
Algorithm 1 DReSS algorithm.
  Input: selected data $X$, number of layers in LLMs $l$, initial model $W$, number of regularization iterations $n$, norm type: flag.
  if flag==1 then
     use Proposition 3.2 to transform problem
  end if
  $X_{1},X_{2}\leftarrow divide(X)$
  for $i=1$ to $n$ do
[…]
     ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$00
     for $j=1$ to $l$ do
        ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{att}}$+$regularize(W_{Q}^{j},W_{K}^{j},W_{V}^{j},W_{o}^{j},X_{1})$
        ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{FFN}}$+$regularize(W_{up}^{j},W_{down}^{j},X_{1})$
        ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{2}}$+$regularize(w_{N}^{j},X_{1})$
     end for
[…]
     ~$\tilde{\mathcal{L}}_{\text{remain}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{1}}$+~$\tilde{\mathcal{L}}_{\text{2}}$+~$\tilde{\mathcal{L}}_{\text{3}}$
     $\mathcal{L}(W,X)\leftarrow forward(W,X_{1})$
     $\mathcal{L}_{\text{sum}}$$\leftarrow$$\mathcal{L}(W,X)$+~$\tilde{\mathcal{L}}_{\text{att}}$+~$\tilde{\mathcal{L}}_{\text{FFN}}$+~$\tilde{\mathcal{L}}_{\text{remain}}$
     update $W$ using BP algorithm
  end for
  $W\leftarrow Prune(W,R)$
  $W\leftarrow Optional\_RFT(W,X_{2})$
3.1 Data Selection
The required data is selected from widely used calibration datasets, including Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020). For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for both pre-pruning regularization and post-pruning RFT. In Section 4.5, we analyze the impact of different datasets.
3.2 Regularization
To provide a clearer understanding of this section, we first introduce the relevant notations. Let $p$ denote the pruning ratio, which indicates that $p\%$ of the model’s parameters need to be pruned. The model dimension of the LLMs is represented by $d$, and the number of layers is denoted by $l$. Additionally, let $W_{emb}\in\mathbb{R}^{b\times n\times d}$ represents the data feature map, and $W_{pos}\in\mathbb{R}^{b\times n\times d}$ denotes the positional feature map, where $b$ is the batch size and $n$ is the number of tokens. The model weights are represented by $W\in\mathbb{R}^{d_{1}\times d_{2}}$. Specifically, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{K}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}^{i}\in\mathbb{R}^{d\times d_{1}}$, and $W_{o}^{i}\in\mathbb{R}^{d_{1}\times d}$ represent the parameter matrices in the $i$th layer Attention block. Similarly, $W_{up}^{i}\in\mathbb{R}^{d\times d_{2}}$ and $W_{down}^{i}\in\mathbb{R}^{d_{2}\times d}$ represent the parameter matrices in the $i$th layer FFN block. Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$-th layer, and $W_{lm}\in\mathbb{R}^{b\times n}$ represents the language modeling matrix. Finally, $I\in\mathbb{R}^{d\times d}$ represents the identity matrix, let $R\in\mathbb{R}^{d\times d}$ represent a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
Figure 2: The pruning strategy of DReSS in a transformer layer.
Proposition 3.1.
(Dependency between the Attention Block and FFN Block.) If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block. The same dependency applies to $W_{o}^{i}$ and $W_{up}^{i-1}$.
We illustrate the process of applying regularization within a transformer layer (Vaswani, 2017), as shown in Figure 2. Following Proposition 3.1, in the Attention block, regularization is applied to the rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ corresponding to the one entries in the index matrix $R$, as well as to the columns of $W_{o}^{i}$ associated with these entries. Similarly, in the FFN block, regularization is enforced on the rows of $W_{up}^{i}$ and the columns of $W_{down}^{i}$ that correspond to the one entries in $R$. We also apply regularization to the remaining parts in LLMs. Specifically, for $W_{emb}$, regularization is enforced on the columns corresponding to the one indices in $R$, and the same procedure is applied to $W_{pos}$. For each $w_{N}^{i}$ within every transformer layer, regularization is applied to the elements corresponding to the one indices in $R$. Additionally, for the final $W_{lm}$, due to its dependency on $W_{down}^{l}$ in the FFN block of the last transformer layer, regularization is applied to the corresponding rows. Therefore, the overall loss function consists of the language modeling loss and the loss induced by regularization, which is defined below:
$\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(W,X)+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}\end{split}$ (1)
where $\mathcal{L}(W,X)$ denotes the language modeling loss and $X$ is the input data, and the remaining term represents the regularization loss. ~$\tilde{\mathcal{L}}_{i}^{\text{att}}$ denotes the $i$th layer attention block regularization loss, which is defined as:
~$\begin{split}\tilde{\mathcal{L}}_{i}^{\text{att}}=\lambda(\|RW_{Q}^{i}\|+\|RW_{K}^{i}\|+\|RW_{V}^{i}\|+\|W_{o}^{i}R\|)\end{split}$ (2)
~$\tilde{\mathcal{L}}_{i}^{\text{FFN}}$ denotes the $i$th layer FFN block regularization loss and the definition is as below:
~$\begin{split}\tilde{\mathcal{L}}_{\text{FFN}}^{i}=\lambda\|RW_{up}^{i}\|+\lambda\|W_{down}^{i}R\|\end{split}$ (3)
~$\tilde{\mathcal{L}}_{\text{remain}}$ denotes the remaining parts of the LLMs’ regularization loss as follows:
~$\begin{split}\tilde{\mathcal{L}}_{\text{remain}}=\lambda(\|W_{emb}R\|+\|W_{pos}R\|+\\ \sum_{i=1}^{l}\|Rw_{N}^{i}\|+\|RW_{lm}\|)\end{split}$ (4)
[…]
Proposition 3.2.
(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation.) The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $|\cdot|_{1}$ denotes the $\ell_{1}$-norm.
$\min\quad|x|_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (5)
s.t. $\quad-y\leq x\leq y,$
$\quad y\geq 0.$
We formalize the minimization of the overall loss as an optimization problem. When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation. In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example. According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by backpropagation.
$\min_{W}\quad\mathcal{L}(W,X)+\lambda|RW_{up}^{i}|_{1}+\lambda|W_{down}^{i}R|_{1}$ (6)
$\min_{W,Y_{1},Y_{2}}$ $\quad\mathcal{L}(W,X)+\mathbf{1}^{T}Y_{1}\mathbf{1}+\mathbf{1}^{T}Y_{2}\mathbf{1}$ (7)
s.t. $\quad-Y_{1}\leq RW_{up}^{i}\leq Y_{1},$
$\quad-Y_{2}\leq W_{down}^{i}R\leq Y_{2},$
$\quad Y_{1}\geq 0,\quad Y_{2}\geq 0.$
During the forward and backward propagation processes, due to the application of regularization on certain specific channels in LLMs, the values within these rows or columns are significantly reduced. Intuitively, this reduction could potentially decrease the impact of these parameters on the model’s performance.
3.3 Pruning
After completing the regularization, we prune the rows and columns directly using the pseudo-index selection matrix $R$. For clarity, we define $S=I-R$ and the pruning for the $i$th layer attention block is then performed as:
${W_{Q}^{i}}^{\prime}$ $=SW_{Q}^{i},\quad{W_{K}^{i}}^{\prime}=SW_{K}^{i}$ (8)
${W_{V}^{i}}^{\prime}$ $=SW_{V}^{i},\quad{W_{o}^{i}}^{\prime}=W_{o}^{i}S$
While for the FFN block, it follows:
${W_{up}^{i}}^{\prime}=SW_{up}^{i},\quad{W_{down}^{i}}^{\prime}=W_{down}^{i}S$ (9)
The remaining parts are pruned as below:
${W_{emb}}^{\prime}$ $=W_{emb}S,\quad{W_{pos}}^{\prime}=W_{pos}S$ (10)
${w_{N}^{i}}^{\prime}$ $=S{w_{N}^{i}},\quad{W_{lm}}^{\prime}=SW_{lm}$
The entire pruning process is straightforward and simple.
3.4 Optional RFT
After pruning, we perform RFT on the model using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al., 2021). This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance. in terms of both perplexity and zero-shot accuracy.
Table 1: Performance comparison of different pruning methods. ’PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 LLM Surgeon 25% 7.26 67.28 63.25 54.24 51.62 35.85 54.44 SliceGPT 25% 7.08 69.21 65.35 52.40 53.70 31.66 54.46 SLEB 25% 7.82 67.94 62.79 49.80 48.55 29.34 51.68 DReSS-$\ell_{2}$ 25% 6.25 68.52 66.71 56.73 52.78 37.63 56.47 DReSS-$\ell_{1}$ 25% 6.28 68.67 65.32 57.16 52.39 37.28 56.16 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 LLM Surgeon 25% 12.60 68.22 58.06 46.25 51.52 27.31 50.27 SliceGPT 25% 14.56 65.29 57.22 47.85 49.79 27.99 49.63 SLEB 25% 16.73 63.28 55.37 43.09 45.62 24.21 46.31 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 LLM Surgeon 25% 11.33 72.12 61.08 62.36 54.19 30.52 56.05 SliceGPT 25% 11.90 70.35 60.62 58.15 52.78 29.52 54.28 SLEB 25% 13.80 71.28 60.96 59.26 51.54 29.83 54.57 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 LLM Surgeon 25% 11.02 74.18 64.33 65.37 60.96 34.98 59.96 SliceGPT 25% 10.94 73.67 64.25 63.28 60.52 34.64 59.27 SLEB 25% 12.02 72.62 63.96 62.79 59.21 34.12 58.50 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 LLM Surgeon 25% 7.38 70.59 65.87 58.66 63.65 38.33 59.42 SliceGPT 25% 7.55 66.87 63.38 54.16 58.46 34.56 55.48 SLEB 25% 10.24 63.25 62.36 53.77 55.82 32.24 53.49 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 LLM Surgeon 25% 5.75 77.75 69.62 74.31 72.83 43.52 67.61 SliceGPT 25% 6.30 68.55 67.48 58.10 62.50 37.88 58.90 SLEB 25% 7.39 66.93 65.87 55.48 60.06 35.14 56.70 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
Figure 3: The ratio of the sum of absolute parameter values after regularization to that before regularization for LLaMA2-7B. For example, regularization is applied to the last 25% of rows in $W_{Q}$, and the absolute sum of these rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Preserved Ratios. This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the regularized columns is compared to the original sum.
4 Experiments
This section introduce experimental setup 4.1 and analyzes the effectiveness of DReSS from the following aspects: performance 4.2, acceleration effectiveness 4.3, robustness under different pruning ratios 4.4, robustness under different datasets 4.5, low overhead 4.6, and ablation study 4.7.
4.1 Experimental Setup
Implementation DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019), utilizing the Hugging Face Transformers library (Wolf, 2019). All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness (Gao et al., 2024a) to evaluate the pruned models.
Datasets To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks. For generation task, following prior work(Ashkboos et al., 2024), we evaluate the model’s perplexity on WikiText-2. For zero-shot task evaluation, the benchmarks consists of PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018). To demonstrate that DReSS does not have a specific dependency on the regularization data, we use a random selection of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration prior to pruning and for the RFT after pruning.
Models The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
Baselines We evaluate DReSS against structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024). Among them, LLM Surgeon and SliceGPT are channel-wise, SLEB is layer-wise. To ensure a fair comparison, in the subsequent experiments, we ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB before pruning, and that the data used for RFT in each method is also the same.
Evaluation Metrics The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024). The acceleration effects are represented by throughput and latency (Song et al., 2024).
4.2 Performance Comparison
To ensure fairness, we randomly selected 1,000 samples from WikiText-2 for each method with a sequence length of 2048. The ratio of calibartion data and the RFT data was set to 3:1, and the pruning ratio was 25%.
As shown in Table 1, DReSS demonstrates superior performance across both generation and zero-shot tasks achieving lowest perplexity and highest average accuracy across all models. For instance, on LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon. Notably, on OPT-13B, compared to the dense model, DReSS experiences only a 1% drop in average accuracy. The performance difference between the $\ell_{2}$-norm and $\ell_{1}$-norm is tiny. In Figure 3, we present the ratio of the absolute values of model parameters and activation values after and before regularization of the selected portions. The results show that, in most intermediate layers, the parameters and activation values after regularization are reduced to 30% of their original values, which indicates that the amount of information contained in this part of the parameter space has decreased. Overall, regularization may help transfer important information hidden in the pruned portions to the remaining parts.
4.3 Acceleration Comparison
Language processing in LLMs comprises two primary stages: prompt processing, which is compute-bound, and token generation, which is memory-bound. We separately analyze the speedup achieved in each stage. Table 2 presents the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated using a single NVIDIA A100 GPU. Following the methodology of previous work (Song et al., 2024), the token generation test scenario involves generating sentences of 128 tokens with a batch size of 64, whereas for prompt processing, latency is measured by processing an input sequence of 2048 tokens.
At a pruning ratio of 50% on OPT-13B, DReSS delivers a 35% improvement in throughput and a 30% reduction in latency compared to the dense model. These results highlight the superior efficiency of DReSS in accelerating model inference.
Table 2: Comparison of throughput and latency under different ratios on OPT-13B and LLaMA2-13B. ‘PPL’ refers to the perplexity on Wikitext2, ‘PR’ represents ‘pruning ratio’, ‘TI’ represents ‘throughput increase’.
Model Method PR PPL($\downarrow$) Tokens/s TI Latency(ms) Speedup OPT-13B Dense 0% 10.12 1029 1.00× 386.5 1.00× DReSS 25% 10.38 1194 1.16× 319.42 1.21× DReSS 50% 12.15 1389 1.35× 274.11 1.41× LLaMA2-13B Dense 0% 4.88 1066 1.00× 396.9 1.00× DReSS 25% 5.12 1215 1.14× 330.8 1.20× DReSS 50% 7.59 1407 1.32× 285.5 1.39×
4.4 Robustness to Varying Pruning Ratios.
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B on WikiText-2 under different methods are shown in Figure 4. DReSS significantly outperforms other methods across various pruning ratios. At a pruning ratio of 20%, DReSS performs similarly to the dense model. When the pruning ratio is up to 60%, SLEB collapses, while DReSS maintains relatively low perplexity compared to other pruning methods. This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to achieve high sparsity levels and unlocking significant potential for model acceleration.
Figure 4: Perplexity of LLaMA-7B pruned by various approaches under different pruning ratios on WikiText-2.
4.5 Dependency on Datasets
Since DReSS relies on data-driven regularization, we investigate its potential dataset dependence. We evaluated perplexity for four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4. To ensure fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2. As shown in Figure 5, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness and scalability.
Figure 5: Comparison of perplexity on different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
4.6 Minimal Overhead
We evaluated the perplexity of each method on LLaMA2-7B using varying amounts of data, keeping all other conditions consistent with Section 4.2 and only change the data size from 500 to 8,000. As shown in Table 3, when only 1,000 samples were used for regularization and RFT, DReSS outperformed the other methods that used 4,000 samples. This further highlights the effectiveness and efficiency of applying regularization prior to pruning, as it transfers critical information from the pruned segments in advance, enabling DReSS to deliver strong performance with minimal data overhead.
Table 3: The impact of the amount of data on model performance.
Data Size 500 1000 2000 4000 8000 LLM Surgeon 8.14 7.38 6.55 5.89 5.56 SliceGPT 8.63 7.55 6.82 6.07 5.74 SLEB 13.26 10.24 9.67 8.79 7.83 DReSS 6.75 5.86 5.61 5.48 5.43
4.7 Ablation Study
Effectiveness of Regularization As shown in Table 4, both random pruning and full-parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio for LLaMA2-7B and LLaMA2-13B. This underscores the importance of applying regularization only to the pruned components. Additionally, RFT after pruning has minimal impact on perplexity and average accuracy, indicating that the RFT process is optional.
Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. “w/o” denotes variants with specific components removed.
Model Method Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B DReSS 5.86 0.00 61.54 0.00 w FPR 12.68 +6.82 50.26 -11.28 w/o R 22.38 +16.52 47.25 -14.29 w/o RFT 5.97 + 0.11 59.56 -1.98 LLaMA2-13B DReSS 5.12 0.00 68.84 0.00 w FPR 10.53 +5.41 53.28 -15.56 w/o R 18.94 +13.82 50.17 -18.67 w/o RFT 5.39 + 0.27 67.96 -0.88
Table 5: The perplexity on WikiText-2 under different regularization and RFT data ratio using LLaMA2-7B at a pruning ratio of 25%.
Ratio (R/F) 4:1 3:1 2:1 1:1 1:2 1:3 PPL($\downarrow$) 5.89 5.86 6.07 6.73 7.45 8.62
Table 6: The perplexity of different $\lambda$ on LLaMA2-7B.
$\lambda$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ $5\times 10^{-2}$ $10^{-1}$ PPL($\downarrow$) 7.47 6.04 5.86 5.93 6.12 7.26 10.55
Data Ratio of Regualrization and RFT With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5. It can be observed that when the ratio is 3:1, the model achieves the lowest perplexity. When the ratio is 4:1, the increase in perplexity is less pronounced, but when the ratio is 1:1, the perplexity increases significantly by 15%. These results underscore the importance of the regularization process relative to the RFT applied after pruning.
Trade Off Between Language Modeling and Regularization Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. As shown in Table 6, when $\lambda=10^{-5}$, the regularization loss is too small to effectively shift important values, leading to a significant increase in perplexity after pruning. In contrast, when $\lambda=10^{-1}$, excessive regularization compromises the model’s language modeling ability, causing a notable performance decline post-pruning. Optimal performance is achieved at $\lambda=10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss for effective pruning.
5 Conclusion
In this paper, we introduce a novel pruning paradigm: first applying regularization to the pruned components, followed by pruning and then RFT. Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during the regularization, effectively mitigating the irreversible performance degradation caused by information loss. DReSS demonstrates superior performance in both generation and zero-shot tasks, significantly outperforming existing pruning methods. For instance, DReSS surpasses the powerful LLM Surgeon by 20% in perplexity on LLaMA2-7B. On OPT-13B, with a pruning ratio of 25%, the average accuracy drops by only 1%, while achieving a 1.41× speedup compared to the dense model. Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing both data and computational costs while minimizing the reliance on RFT.
Impact Statement
Pruning plays a crucial role in model compression. DReSS is an efficient and low-overhead pruning method, which enables smaller models to achieve competitive performance, especially in resource-constrained scenarios and it also provides some insights for the development of LLMs.

### v2

@@
 Abstract
-Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale results in high computational and memory costs.
-Recent studies have revealed that LLMs exhibit sparsity, providing the potential to reduce model size through pruning techniques.
+Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale leads to high computational and memory costs.
+Recent studies show that LLMs exhibit sparsity, which can be exploited for pruning.
 However, existing pruning methods typically follow a prune-then-finetune paradigm.
-Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, imposing a substantial computational burden to recover performance during finetuning.
-In this paper, we propose a novel paradigm that first applies regularization, then prunes, and finally finetunes.
+Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, necessitating expensive fine-tuning to recover performance.
+To address this, we propose a new paradigm: first apply regularization, then prune, and finally fine-tune.
 Based on this paradigm, we introduce DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs.
-By leveraging a small amount of data to regularize the components to be pruned, DReSS explicitly transfers the important information to the remaining parts of the model in advance.
+By leveraging a small amount of data to regularize the components before pruning, DReSS transfers the important information to the remaining parts of the model in advance.
 Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities.
-Experimental results demonstrate that DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
+Experimental results demonstrate DReSS achieves comparable performance to conventional methods even without fine-tuning, drastically alleviating computational costs.
+Moreover, DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
 Machine Learning, ICML
@@
 Large language models (LLMs) have achieved significant advancements across a wide range of tasks and domains, demonstrating their robust capabilities (Zhang et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Wu et al., 2024).
-However as the model size grows, the increased number of parameters results in substantial computational and memory demands.
-These factors severely hinder the deployment of LLMs in practical applications.
-Therefore, it is urgent to develop methods that can reduce model size while maintaining performance.
+However, as the model size increases, the growing number of parameters leads to significant computational and memory requirements, which significantly hinder the practical deployment of LLMs.
+Consequently, it is urgent to develop methods that can reduce model size while maintaining performance.
 To address these challenges, several methods have been proposed, including pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Ma et al., 2023; An et al., 2024), quantization (Frantar et al., 2022; Xiao et al., 2023), knowledge distillation (Shridhar et al., 2022; Hsieh et al., 2023), and low-rank decomposition (Saha et al., 2023).
-In this work, we mainly focus on pruning, an efficient and highly generalizable method that can be combined with other model compression strategies.
+In this work, we mainly focus on pruning—an efficient and highly generalizable approach that can be seamlessly integrated with other model compression strategies.
 Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and structured pruning (Ma et al., 2023; An et al., 2024).
-Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) on the pruned model for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024).
-Moreover, the model obtained through structured pruning typically achieves much faster inference speed, while ensuring the same perplexity and accuracy as the model obtained through unstructured pruning.
+Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024).
+Moreover, the model obtained through structured pruning typically achieves much faster inference speed due to the regular data patterns.
 Despite these advancements, existing structured pruning methods still have some limitations.
-They all follow the paradigm of first selecting channels or layers for pruning based on a designed metric, and then performing RFT (Chavan et al., 2024).
-However, they neglect the fact that important information can also exist in the pruned channels or layers (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance.
-Thus, in order to achieve satisfactory performance with the pruned model, a large amount of data must be introduced for RFT (Ma et al., 2023), which increases computational overhead.
-Moreover, they often lead to a collapse in model performance at high pruning ratios, limiting their effectiveness in reducing lantency and enhancing throughput.
+They all follow the paradigm of first selecting channels or layers to prune based on a designed metric, and then performing RFT (Chavan et al., 2024).
+However, they neglect the fact that important information can also exist in the pruned parts (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance.
+Thus, achieving satisfactory performance with pruned models often requires extensive data for RFT (Ma et al., 2023), increasing computational costs.
+Additionally, high pruning ratios frequently cause performance collapse, limiting their effectiveness in reducing latency and improving throughput.
 To tackle these limitations, in this paper, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then removes those parameters, and finally performs RFT.
@@
 The comparison of the proposed method with existing channel-wise pruning approaches is shown in Figure 1.
-Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, then (4) do RFT.
-First, we randomly selecting a small subset of data from widely used benchmark datasets, as the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced.
+Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, finally (4) do RFT.
+First, we randomly select a small subset of data from widely used benchmark datasets.
+Since the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced.
 Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization to these channels using the selected data.
-This shifts important information to the unpruned parts, significantly mitigating performance degradation caused by direct pruning and ensuring the model maintains strong performance even at high pruning ratios, thus enabling substantial inference speedup.
+This redistributes important information to the unpruned components, thereby significantly mitigating performance degradation caused by direct pruning.
+As a result, the model can maintain robust performance even at high pruning ratios, facilitating a substantial speedup in inference.
 Third, we prune the parameters that have been regularized.
-Finally, we do RFT on the pruned model using a small subset of samples selected in the first step.
-Extensive experiments demonstrate that, DReSS significantly surpasses existing pruning methods in terms of perplexity and accuracy across various pruning ratios, it also substantially improves inference speed.
+Finally, we perform RFT on the pruned model using a small subset of samples selected in the first step.
+Extensive experiments show that DReSS significantly outperforms existing pruning methods in perplexity and accuracy across various pruning ratios, even without RFT, while also greatly enhancing inference speed.
 The main contributions of our work are summarized as follows:
-  • Propose A New Paradigm: By first applying regularization, then pruning, and finally performing RFT, DReSS reduces information loss from direct parameter removal, enhancing the model’s language modeling ability and lowering the computational cost of RFT.
-  • High Performance: DReSS outperforms existing structured pruning methods on both generation and zero-shot tasks.
-The pruned model shows a significant increase in throughput and a notable reduction in latency compared to the dense model.
-  • Low Overhead: By using only a small amount of data during the regularization and optional RFT process, the pruning overhead is minimal.
+  • Propose A New Paradigm: By sequentially applying regularization, pruning, and RFT, DReSS minimizes information loss caused by direct parameter removal, thereby improving the model’s language modeling capabilities and reducing the computational cost of RFT.
+  • High Performance: DReSS surpasses existing structured pruning methods in both generation and zero-shot tasks.
+The pruned model achieves a substantial increase in throughput and a significant reduction in latency compared to the dense model.
+  • Low Overhead: By leveraging only a small amount of data for regularization and optional RFT, DReSS incurs minimal overhead.
 Figure 1: A comparison between existing channel-wise structured pruning methods and DReSS.
@@
 The taller orange-yellow cylinder represents larger data volume.
-Here, L($\cdot$) represents the regularization, and S represents pruning.
+L($\cdot$) represents the regularization, and S represents pruning.
 Above: Existing methods first select channels based on importance, followed by pruning and RFT on the pruned model.
-Below: DReSS first regularizes channels to minimize performance impact, prunes the channels, then do RFT with a little amount of data.
+Below: DReSS first regularizes channels to transfer important information, prunes the channels, then do RFT with a little amount of data.
 2 Related Works
@@
 Unstructured pruning methods remove individual weights based on their importance.
-For instance, Magnitude pruning (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix.
+For instance, Magnitude (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix.
 The main advantage is flexibility, enabling the removal of unimportant weights while retaining critical ones, allowing high sparsity in LLMs.
-However, its limitations include the need for specialized hardware for accelerated inference (Xia et al., 2023a), high sparsity for significant acceleration (Wang, 2020), and inability to perform RFT on downstream tasks.
+However, its limitations include the requirement for specialized hardware to accelerate inference (Xia et al., 2023a), the necessity for high sparsity to achieve substantial acceleration (Wang, 2020), and the inability to perform RFT on downstream tasks.
 Structured pruning methods, including channel-wise pruning (An et al., 2024) and layer-wise pruning (Men et al., 2024), can achieve significant inference speedup with a low sparsity.
@@
 Both of them can significantly alter the distribution pattern of the data (Han et al., 2015; Tao et al., 2023).
-Motivated by this observation, we propose a data-driven regularization approach for parameters designated for pruning.
-By applying a small amount of data, the regularization probably transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
+Motivated by this observation, we propose a data-driven regularization approach.
+By applying a small amount of data, the regularization transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
 3 Methodology
@@
   • Data Selection: Select a small amount of data for pre-pruning regularization and post-pruning RFT.
-  • Regularization: Applying regularization to the selected channels of weight matrices during both forward and backward propagations.
+  • Regularization: Applying regularization to the selected channels of weight matrices during forward and backward propagations.
   • Pruing: Prune the regularized portions.
@@
      $\mathcal{L}_{\text{sum}}$$\leftarrow$$\mathcal{L}(W,X)$+~$\tilde{\mathcal{L}}_{\text{att}}$+~$\tilde{\mathcal{L}}_{\text{FFN}}$+~$\tilde{\mathcal{L}}_{\text{remain}}$
-     update $W$ using BP algorithm
+     update $W$ using backpropagation algorithm
   end for
@@
 The required data is selected from widely used calibration datasets, including Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020).
-For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for both pre-pruning regularization and post-pruning RFT.
-In Section 4.5, we analyze the impact of different datasets.
+For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for pre-pruning regularization and post-pruning RFT.
+In Section 4.5, we analyze the impact of using different datasets.
 3.2 Regularization
@@
 Similarly, $W_{up}^{i}\in\mathbb{R}^{d\times d_{2}}$ and $W_{down}^{i}\in\mathbb{R}^{d_{2}\times d}$ represent the parameter matrices in the $i$th layer FFN block.
-Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$-th layer, and $W_{lm}\in\mathbb{R}^{b\times n}$ represents the language modeling matrix.
+Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$th layer, and $W_{lm}\in\mathbb{R}^{d\times b\times n}$ represents the language modeling matrix.
 Finally, $I\in\mathbb{R}^{d\times d}$ represents the identity matrix, let $R\in\mathbb{R}^{d\times d}$ represent a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
@@
 Proposition 3.1.
-(Dependency between the Attention Block and FFN Block.) If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block.
+(Dependency between the Attention Block and FFN Block.
+Proof in Appendix A.1).
+If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block.
 The same dependency applies to $W_{o}^{i}$ and $W_{up}^{i-1}$.
@@
 Proposition 3.2.
-(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation.) The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $|\cdot|_{1}$ denotes the $\ell_{1}$-norm.
-$\min\quad|x|_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (5)
+(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation.
+Proof in Appendix A.2).
+The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
+$\min\quad||x||_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (5)
 s.t. $\quad-y\leq x\leq y,$
@@
 We formalize the minimization of the overall loss as an optimization problem.
-When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation.
+When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation (BP) algorithm (Rumelhart et al., 1986).
 In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example.
-According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by backpropagation.
-$\min_{W}\quad\mathcal{L}(W,X)+\lambda|RW_{up}^{i}|_{1}+\lambda|W_{down}^{i}R|_{1}$ (6)
+According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by BP algorithm.
+$\min_{W}\quad\mathcal{L}(W,X)+\lambda||RW_{up}^{i}||_{1}+\lambda||W_{down}^{i}R||_{1}$ (6)
 $\min_{W,Y_{1},Y_{2}}$ $\quad\mathcal{L}(W,X)+\mathbf{1}^{T}Y_{1}\mathbf{1}+\mathbf{1}^{T}Y_{2}\mathbf{1}$ (7)
@@
 During the forward and backward propagation processes, due to the application of regularization on certain specific channels in LLMs, the values within these rows or columns are significantly reduced.
-Intuitively, this reduction could potentially decrease the impact of these parameters on the model’s performance.
+Intuitively, this reduction could potentially transfer important information and decrease the impact of these parameters on the model’s performance.
 3.3 Pruning
@@
 After pruning, we perform RFT on the model using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al., 2021).
-This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance.
-in terms of both perplexity and zero-shot accuracy.
+This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance in terms of both perplexity and zero-shot accuracy.
 Table 1: Performance comparison of different pruning methods.
-’PPL’ refers to the perplexity on WikiText-2.
+‘PPL’ refers to the perplexity on WikiText-2.
 The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy.
@@
 4 Experiments
-This section introduce experimental setup 4.1 and analyzes the effectiveness of DReSS from the following aspects: performance 4.2, acceleration effectiveness 4.3, robustness under different pruning ratios 4.4, robustness under different datasets 4.5, low overhead 4.6, and ablation study 4.7.
+This section introduces experimental setup (4.1) and analyzes the effectiveness of DReSS from the following aspects: performance comparison (4.2), acceleration (4.3), robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (4.7).
 4.1 Experimental Setup
-Implementation DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019), utilizing the Hugging Face Transformers library (Wolf, 2019).
+Implementation: DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019) , utilizing the Hugging Face Transformers library (Wolf, 2019).
 All experiments are conducted on 80GB NVIDIA A100 GPUs.
 For fairness, we use llm-eval-harness (Gao et al., 2024a) to evaluate the pruned models.
-Datasets To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks.
+More details are shown in Appendix B.
+Datasets: To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks.
 For generation task, following prior work(Ashkboos et al., 2024), we evaluate the model’s perplexity on WikiText-2.
 For zero-shot task evaluation, the benchmarks consists of PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018).
-To demonstrate that DReSS does not have a specific dependency on the regularization data, we use a random selection of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration prior to pruning and for the RFT after pruning.
-Models The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
-Baselines We evaluate DReSS against structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024).
+To demonstrate that DReSS is not specifically dependent on the regularization data, we use a randomly selected subset of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration before pruning and for RFT after pruning.
+Models: The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
+Baselines: We evaluate DReSS against competitive structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024).
 Among them, LLM Surgeon and SliceGPT are channel-wise, SLEB is layer-wise.
 To ensure a fair comparison, in the subsequent experiments, we ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB before pruning, and that the data used for RFT in each method is also the same.
-Evaluation Metrics The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024).
+We also compared with powerful unstructured pruning methods like Mangniude (Han et al., 2015), Wanda (Sun et al., 2023), and SparseGPT (Frantar & Alistarh, 2023) and the results are in Appendix C.
+Evaluation Metrics: The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024).
 The acceleration effects are represented by throughput and latency (Song et al., 2024).
@@
 The results show that, in most intermediate layers, the parameters and activation values after regularization are reduced to 30% of their original values, which indicates that the amount of information contained in this part of the parameter space has decreased.
-Overall, regularization may help transfer important information hidden in the pruned portions to the remaining parts.
-4.3 Acceleration Comparison
+We also present the changes occurring in the unregularized part in Appendix D.
+The results show that this part has increased compared to before, indicating that it now accommodates more important information.
+Overall, regularization transfers important information hidden in the pruned portions to the remaining parts.
+4.3 Acceleration Effectiveness
 Language processing in LLMs comprises two primary stages: prompt processing, which is compute-bound, and token generation, which is memory-bound.
@@
 Model Method PR PPL($\downarrow$) Tokens/s TI Latency(ms) Speedup OPT-13B Dense 0% 10.12 1029 1.00× 386.5 1.00× DReSS 25% 10.38 1194 1.16× 319.42 1.21× DReSS 50% 12.15 1389 1.35× 274.11 1.41× LLaMA2-13B Dense 0% 4.88 1066 1.00× 396.9 1.00× DReSS 25% 5.12 1215 1.14× 330.8 1.20× DReSS 50% 7.59 1407 1.32× 285.5 1.39×
-4.4 Robustness to Varying Pruning Ratios.
+4.4 Robustness to Varying Pruning Ratios
 Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%.
@@
 This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to achieve high sparsity levels and unlocking significant potential for model acceleration.
+A more detailed exploration of pruning ratios using DReSS is provided in the Appendix E.
 Figure 4: Perplexity of LLaMA-7B pruned by various approaches under different pruning ratios on WikiText-2.
@@
 As shown in Figure 5, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness and scalability.
+More details are shown in Appendix F.
 Figure 5: Comparison of perplexity on different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
@@
 4.7 Ablation Study
-Effectiveness of Regularization As shown in Table 4, both random pruning and full-parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio for LLaMA2-7B and LLaMA2-13B.
-This underscores the importance of applying regularization only to the pruned components.
-Additionally, RFT after pruning has minimal impact on perplexity and average accuracy, indicating that the RFT process is optional.
-Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. “w/o” denotes variants with specific components removed.
+Effectiveness of Regularization: As shown in Table 4, both random pruning and full parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio.
+This highlights the importance of applying regularization exclusively to the pruned components.
+Additionally, RFT after pruning has minimal impact on model performance, indicating that the RFT process is optional.
+Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. ‘w/o’ denotes the removal of specific components.
 Model Method Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B DReSS 5.86 0.00 61.54 0.00 w FPR 12.68 +6.82 50.26 -11.28 w/o R 22.38 +16.52 47.25 -14.29 w/o RFT 5.97 + 0.11 59.56 -1.98 LLaMA2-13B DReSS 5.12 0.00 68.84 0.00 w FPR 10.53 +5.41 53.28 -15.56 w/o R 18.94 +13.82 50.17 -18.67 w/o RFT 5.39 + 0.27 67.96 -0.88
@@
 $\lambda$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ $5\times 10^{-2}$ $10^{-1}$ PPL($\downarrow$) 7.47 6.04 5.86 5.93 6.12 7.26 10.55
-Data Ratio of Regualrization and RFT With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5.
+Data Ratio of Regualrization and RFT: With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5.
 It can be observed that when the ratio is 3:1, the model achieves the lowest perplexity.
@@
 These results underscore the importance of the regularization process relative to the RFT applied after pruning.
-Trade Off Between Language Modeling and Regularization Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$.
+Trade-Off Between Language Modeling and Regularization: Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$.
 As shown in Table 6, when $\lambda=10^{-5}$, the regularization loss is too small to effectively shift important values, leading to a significant increase in perplexity after pruning.
@@
 In this paper, we introduce a novel pruning paradigm: first applying regularization to the pruned components, followed by pruning and then RFT.
-Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during the regularization, effectively mitigating the irreversible performance degradation caused by information loss.
+Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during regularization, effectively mitigating the irreversible performance degradation caused by information loss.
 DReSS demonstrates superior performance in both generation and zero-shot tasks, significantly outperforming existing pruning methods.
@@
 DReSS is an efficient and low-overhead pruning method, which enables smaller models to achieve competitive performance, especially in resource-constrained scenarios and it also provides some insights for the development of LLMs.
+Appendix A Proofs of the Propositions
+A.1 Proof of Proposition 3.1
+Let us first assume that $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times k}$.
+We can express $A$ as $\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}$, where $\mathbf{a}_{i}\in\mathbb{R}^{m\times 1}$, and $B$ as $\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}$, where $\mathbf{b}_{i}\in\mathbb{R}^{k\times 1}$.
+Then, we obtain $C\in\mathbb{R}^{m\times k}$ as shown in Equation 11.
+It is important to note that the final summation in Equation 11 refers to the summation of $n$ rank-1 matrices, not a scalar summation.
+$\begin{split}C=AB=\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}\times\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}=\sum_{i=1}^{n}\mathbf{a}_{i}\mathbf{b}_{i}^{T}\end{split}$ (11)
+According to Equation 11, when some columns of $A$ are zero, the rows of $B$ that correspond to the outer product with the zero columns of $A$ can take any values, as the product of zero and any value is zero.
+These rows can therefore be set to zero.
+Applying the above conclusion to the parameters of a transformer layer, if certain columns of $W_{down}^{i-1}$ in the FFN layer of the $(i-1)$th transformer layer are pruned, then, since $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th transformer layer are immediately multiplied by the result of $W_{down}^{i-1}$, the corresponding rows in $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ can be set to zero.
+Additionally, if certain columns of $W_{o}^{i}$ in the $i$th transformer layer are pruned, the $W_{up}^{i}$ in the FFN block will immediately multiply with it, so the corresponding rows in $W_{up}^{i}$ should also be pruned.
+Thus, we have completed the proof of Proposition 3.1.
+A.2 Proof of Proposition 3.2
+Step 1: Expressing $\ell_{1}$-norm Using Elements.
+The objective function in the unconstrained problem is the $\ell_{1}$-norm of the vector $x$, which is defined as:
+$||x||_{1}=\sum_{i=1}^{n}|x_{i}|$ (12)
+This function aims to minimize the sum of the absolute values of the components of $x$.
+Step 2: Reformulating the Constrained Problem
+The constrained optimization problem introduces an auxiliary variable $y$, where for each element $i$:
+$x_{i}\geq-y_{i}\quad\text{and}\quad x_{i}\leq y_{i}\quad$ (13)
+This implies that $y_{i}\geq|x_{i}|$, meaning each element of $y$ serves as an upper bound for the absolute value of the corresponding element in $x$.
+Consequently, minimizing $|x|_{1}$ is equivalent to minimizing the sum of the elements in $y$.
+Thus, the objective function is defined as:
+$\mathbf{1}^{T}y$ (14)
+Thus, minimizing $\mathbf{1}^{T}y$ is equivalent to minimizing the sum of the absolute values of $x$, which is the $\ell_{1}$-norm of $x$.
+This transformation allows the optimization problem to be solved without directly involving the absolute value function, resulting in an equivalent constrained optimization problem that can be addressed via backpropagation.
+Thus, the proof of the Proposition 3.2 is complete.
+Appendix B Detailed Implementation
+In this part, we first introduce several hyperparameter settings, with the detailed results shown in Table 7.
+In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, OPT-6.7B, OPT-13B, LLaMA2-7B, and LLaMA2-13B.
+For all RFT configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048.
+All other hyperparameters follow the default settings provided in the Hugging Face PEFT package (Mangrulkar et al., 2022).
+Due to limited computational resources, we set the batch size to 64.
+In future work, we will further explore the impact of a broader range of batch sizes on the performance of the pruning method.
+These choices ensure a balance between model adaptability and computational efficiency, leveraging parameter-efficient fine-tuning while maintaining stable optimization dynamics.
+Table 7: Implementation Details
+Precision LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold Min Delta FP16 32 10 2048 64 2e-5 5 1e-4
+To ensure a fair comparison between DReSS and other methods, we maintain consistency in the data used across all approaches.
+Specifically, the data used by DReSS for regularization, by LLM Surgeon for periodic updates of model weights and structures, by SliceGPT for selecting channel importance, and by SLEB for identifying crucial transformer layers are identical.
+Furthermore, we ensure that the data employed during the RFT process is consistent across all methods, thereby enabling a controlled and equitable evaluation framework.
+Following previous works (Ashkboos et al., 2024; Song et al., 2024), for the comparison unstructured pruning methods like Mangniude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by DReSS for regularization.
+Appendix C Results of Comparison with Unstructured Pruning Methods
+We compare DReSS at a 25% pruning ratio with unstructured pruning methods employing 2:4 sparsity (where two out of every four consecutive parameters are set to zero) and present the experimental results in Table 8.
+For unstructured pruning methods, only calibration data is required.
+To ensure consistency, all methods utilize a total of 1,000 samples, randomly selected from WikiText-2.
+The models used in these experiments consists of the OPT and LLaMA model families.
+Compare to structured pruning methods, the primary advantage of unstructured pruning lies in its flexibility, enabling the removal of specific weights while preserving critical information in the remaining parameters of LLMs.
+Based on the results in Table 8, it can be seen that DReSS with a pruning ratio of 25% significantly outperforms existing unstructured pruning methods with a sparsity of 2:4 on generative and zero-shot tasks, demonstrating the powerful performance of DReSS.
+The reason DReSS outperforms unstructured pruning in comparisons lies in the regularization process before pruning.
+This process transfers important information hidden in the parameter space to the remaining parts of the model, preventing irreversible performance degradation caused by direct pruning and the resulting loss of information.
+Table 8: Performance comparison with unstructured pruning methods.
+‘PPL’ refers to the perplexity on WikiText-2.
+The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy.
+The best result is highlighted in bold, and the second-best in underlined.
+DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
+Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 Magnitude 2:4 265 57.46 42.19 35.62 32.58 16.42 36.85 Wanda 2:4 16.54 64.32 57.84 41.63 43.47 23.44 46.14 SparseGPT 2:4 13.48 68.58 56.72 45.36 46.89 25.92 48.69 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 Magnitude 2:4 969 62.16 48.74 43.48 41.57 20.63 43.32 Wanda 2:4 13.29 71.76 60.22 54.20 51.18 28.33 53.14 SparseGPT 2:4 11.55 74.21 60.77 57.25 53.03 29.44 54.94 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 Magnitude 2:4 1.2e4 64.38 52.19 49.46 44.58 21.25 46.37 Wanda 2:4 13.49 72.25 61.72 57.97 53.96 29.69 55.12 SparseGPT 2:4 11.17 74.16 62.43 59.18 56.27 31.74 56.76 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 Magnitude 2:4 54.59 64.42 60.93 56.72 59.18 27.31 53.71 Wanda 2:4 11.35 70.84 62.27 55.33 57.58 31.91 55.59 SparseGPT 2:4 8.67 72.14 64.96 58.93 60.90 34.22 58.23 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 Magnitude 2:4 8.33 70.47 62.04 50.65 62.46 31.74 55.47 Wanda 2:4 8.32 73.94 67.01 63.09 64.31 37.80 61.23 SparseGPT 2:4 7.07 75.46 68.51 65.52 66.04 39.76 63.06 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
+Appendix D Changes in the Parts Without Regularization
+As illustrated in Figure 6, the magnitude of the unregularized parameters exhibits an increase after the application of regularization, suggesting a redistribution of model capacity.
+This phenomenon indicates that during regularization, critical information, initially encoded in the regularized portion of the model, is partially transferred to the unregularized portion.
+In contrast, Figure 3 shows a reduction in the magnitude of the regularized parameters after regularization, implying that the imposed constraints effectively suppress the corresponding parameter values, enforcing sparsity or compression in that region.
+Collectively, these observations suggest that the regularization process facilitates an implicit redistribution of information across different parameter subsets.
+Specifically, the regularization term promotes a shift of important model characteristics from the constrained (regularized) portion to the unconstrained (unregularized) portion, thereby preserving essential model knowledge despite the imposed sparsity constraints.
+Based on this insight, we posit that pruning the regularized portion post-regularization could mitigate information loss, as the core knowledge has already been migrated to the unregularized segment.
+This pruning strategy effectively reduces parameter redundancy while retaining the model’s language modeling capacity, thereby achieving a more compact and efficient representation without compromising performance.
+Figure 6: Ratio of the sum of absolute values of unregularized parameters after LLaMA2-7B regularization to the sum of absolute values of the corresponding parameters before regularization.
+For example, regularization is not applied to the first 75% of rows in $W_{Q}$, and the absolute sum of the rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Increased Ratios.
+This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the unregularized columns is compared to the original sum.
+Appendix E Performance of DReSS under Different Pruning Ratios and Datasets
+E.1 The Perplexity of DReSS under Different Pruning Ratios and Datasets
+In Section 4.2, we utilize 1,000 samples randomly selected from the WikiText-2 dataset to guide the regularization process.
+Subsequently, we evaluate multiple large language models (LLMs) by measuring changes in perplexity across various generative task datasets, including WikiText-2, Alpaca, PTB, and C4, under pruning rates of 10%, 20%, 30%, 40%, 50%, and 60%.
+The detailed results, presented in Table 9, indicate that DReSS exhibits greater robustness as model scale increases, suggesting that the proposed method effectively mitigates performance degradation in larger architectures.
+This highlights the scalability of DReSS and its potential to maintain model efficiency under varying levels of sparsity.
+Table 9: Perplexity comparison of DReSS with different pruning ratios.
+We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the perplexity of the OPT and LLaMA2 models on the generation task datasets Alpaca, WikiText-2, PTB, and C4.
+For DReSS, we use the $\ell_{2}$-norm.
+Model Pruning Ratio WikiText-2 Alpaca PTB C4 OPT-2.7B Dense 12.46 11.64 17.97 14.32 10% 12.48 11.71 18.16 14.71 20% 12.51 11.93 19.45 15.42 30% 14.49 12.88 23.58 18.93 40% 18.92 15.46 31.30 24.33 50% 24.57 21.37 45.32 32.15 60% 33.83 31.22 58.73 45.56 OPT-6.7B Dense 10.85 10.27 15.77 12.71 10% 10.91 10.45 16.05 13.18 20% 11.02 10.83 17.58 14.45 30% 12.32 11.46 19.65 16.68 40% 14.26 12.71 25.52 20.94 50% 19.63 15.66 33.78 28.22 60% 28.75 21.39 47.29 39.47 OPT-13B Dense 10.12 9.46 14.52 12.06 10% 10.26 9.65 14.88 12.37 20% 10.47 9.97 15.64 13.15 30% 10.99 10.56 18.81 15.63 40% 11.62 11.25 23.07 19.55 50% 16.15 14.68 29.26 26.21 60% 27.63 20.12 38.59 35.58 LLaMA2-7B Dense 5.47 5.25 7.92 7.26 10% 5.54 5.29 8.06 7.34 20% 5.63 5.37 8.29 7.79 30% 7.39 7.32 9.13 8.46 40% 9.62 8.62 12.37 10.56 50% 13.77 12.59 18.85 15.18 60% 24.38 20.25 29.34 27.37 LLaMA2-13B Dense 4.88 4.63 7.16 6.73 10% 4.94 4.69 7.23 6.96 20% 5.08 4.83 7.61 7.53 30% 5.61 5.42 8.33 8.14 40% 6.25 6.14 9.27 9.05 50% 7.59 7.28 11.58 10.88 60% 12.77 11.78 15.16 13.62
+E.2 The Accuracy of DReSS under Different Pruning Ratios on Zero-shot Tasks
+To systematically evaluate the performance of DReSS on zero-shot tasks under varying pruning rates, we adopt the experimental setup outlined in Section 4.2, where 1,000 samples are randomly selected from the WikiText-2 dataset to guide the regularization process.
+We assess the accuracy of different model configurations at pruning rates of 10%, 20%, 30%, 40%, 50%, and 60% across a diverse set of benchmark datasets, including PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c.
+The results, summarized in Table 10, provide insights into the impact of sparsity on zero-shot generalization.
+Notably, the analysis reveals that DReSS maintains competitive performance even at higher pruning rates, demonstrating its effectiveness in preserving reasoning and commonsense understanding across different tasks.
+Table 10: Accuracy comparison of DReSS with different pruning ratios.
+We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the accuracy of the OPT and LLaMA2 models on the zero-shot task datasets PIQA, WinoGrande, HellaSwag, ARC-e and ARC-c.
+For DReSS, we use the $\ell_{2}$-norm.
+‘Avg_Acc’ represents the average accuracy.
+Model Pruning Ratio PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39 10% 70.38 60.12 51.79 52.46 29.78 52.91 20% 69.96 59.47 50.45 51.87 28.62 52.07 30% 64.52 58.76 49.25 50.23 27.52 50.06 40% 61.32 56.27 47.39 48.26 25.57 47.76 50% 59.38 53.46 44.63 46.19 21.66 45.06 60% 52.99 48.74 40.03 43.35 16.25 40.27 OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70 10% 75.89 64.61 64.69 58.53 33.47 59.44 20% 75.12 64.23 62.56 57.34 32.95 58.44 30% 72.52 62.63 58.27 54.48 29.99 55.58 40% 67.37 58.59 52.12 49.38 26.87 50.87 50% 61.48 55.62 46.46 47.68 22.97 46.84 60% 55.73 51.52 43.38 45.85 18.62 43.02 OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79 10% 75.46 64.69 68.56 61.79 35.58 61.22 20% 74.78 64.61 67.25 61.66 35.43 60.75 30% 72.62 63.26 65.69 59.64 32.57 58.76 40% 68.67 61.49 62.74 55.98 29.26 55.63 50% 62.19 56.46 58.55 52.15 24.53 50.78 60% 57.43 52.72 51.23 47.62 21.05 46.01 LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00 10% 77.38 68.16 71.28 69.26 43.73 65.96 20% 76.42 67.35 68.26 66.73 41.68 64.09 30% 72.29 64.87 58.53 63.48 39.27 59.69 40% 68.66 61.49 55.42 58.61 36.52 56.14 50% 61.32 55.68 51.79 52.35 31.55 50.54 60% 56.45 51.79 48.96 48.98 27.26 46.69 LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76 10% 79.35 72.15 76.42 76.92 48.57 70.68 20% 78.27 71.98 73.20 75.43 47.62 69.30 30% 75.49 69.73 71.54 72.87 45.41 67.00 40% 73.56 64.46 67.75 68.45 41.28 63.10 50% 68.73 59.82 62.48 60.12 36.14 57.46 60% 62.25 56.27 59.93 53.37 29.77 52.32
+Appendix F More Details of the Dependency on Calibration Dataset
+When using Alpaca, WikiText-2, C4, and PTB as calibration and RFT data, the perplexity of various methods on Alpaca, C4, and PTB is shown as follows:
+Figure 7: Comparison of perplexity on Alpaca using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
+Figure 8: Comparison of perplexity on C4 using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
+Figure 9: Comparison of perplexity on PTB using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.

### v3

@@
 Recent studies show that LLMs exhibit sparsity, which can be exploited for pruning.
-However, existing pruning methods typically follow a prune-then-finetune paradigm.
-Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, necessitating expensive fine-tuning to recover performance.
+Existing pruning methods typically follow a prune-then-finetune paradigm.
+Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, causing expensive fine-tuning to recover performance.
 To address this, we propose a new paradigm: first apply regularization, then prune, and finally fine-tune.
-Based on this paradigm, we introduce DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs.
-By leveraging a small amount of data to regularize the components before pruning, DReSS transfers the important information to the remaining parts of the model in advance.
+Based on this paradigm, we propose DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs.
+By using a small amount of data to regularize the components before pruning, DReSS transfers the important information to the remaining parts of the model in advance.
 Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities.
-Experimental results demonstrate DReSS achieves comparable performance to conventional methods even without fine-tuning, drastically alleviating computational costs.
-Moreover, DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
-Machine Learning, ICML
+We evaluate our method on various LLMs, including Phi-2, OPT, LLaMA2, LLaMA3.
+Experimental results demonstrate DReSS even without recovery fine-tuning (RFT) achieves comparable performance to previous methods, drastically alleviating computational costs.
+Moreover, DReSS significantly outperforms existing powerful pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
 1 Introduction
-Large language models (LLMs) have achieved significant advancements across a wide range of tasks and domains, demonstrating their robust capabilities (Zhang et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Wu et al., 2024).
+Large language models (LLMs) have achieved significant advancements across a wide range of tasks, demonstrating their robust capabilities (Zhang et al.
+2022; Achiam et al.
+2023; Touvron et al.
+2023; Wu et al.
+2024).
 However, as the model size increases, the growing number of parameters leads to significant computational and memory requirements, which significantly hinder the practical deployment of LLMs.
-Consequently, it is urgent to develop methods that can reduce model size while maintaining performance.
-To address these challenges, several methods have been proposed, including pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Ma et al., 2023; An et al., 2024), quantization (Frantar et al., 2022; Xiao et al., 2023), knowledge distillation (Shridhar et al., 2022; Hsieh et al., 2023), and low-rank decomposition (Saha et al., 2023).
-In this work, we mainly focus on pruning—an efficient and highly generalizable approach that can be seamlessly integrated with other model compression strategies.
-Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and structured pruning (Ma et al., 2023; An et al., 2024).
-Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024).
-Moreover, the model obtained through structured pruning typically achieves much faster inference speed due to the regular data patterns.
-Despite these advancements, existing structured pruning methods still have some limitations.
-They all follow the paradigm of first selecting channels or layers to prune based on a designed metric, and then performing RFT (Chavan et al., 2024).
-However, they neglect the fact that important information can also exist in the pruned parts (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance.
-Thus, achieving satisfactory performance with pruned models often requires extensive data for RFT (Ma et al., 2023), increasing computational costs.
-Additionally, high pruning ratios frequently cause performance collapse, limiting their effectiveness in reducing latency and improving throughput.
-To tackle these limitations, in this paper, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then removes those parameters, and finally performs RFT.
+Consequently, it is critical to develop methods that can reduce model size while maintaining performance.
+To address these challenges, several methods have been proposed, including pruning (Frantar and Alistarh 2023; Sun et al.
+2024; An et al.
+2024), quantization (Frantar et al.
+2023; Xiao et al.
+2023), knowledge distillation (Shridhar, Stolfo, and Sachan 2023; Hsieh et al.
+2023), and low-rank decomposition (Saha, Srivastava, and Pilanci 2023).
+In this work, we mainly focus on pruning which is an efficient and highly generalizable approach that can be seamlessly integrated with other model compression strategies.
+Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar and Alistarh 2023; Sun et al.
+2024) and structured pruning (Ma, Fang, and Wang 2023; An et al.
+2024).
+Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) for specific downstream tasks without relying on specialized hardware (Zhu et al.
+2024).
+The model obtained through structured pruning typically achieves much faster inference speed due to the regular data patterns.
+Despite these advancements, existing structured pruning methods still have limitations.
+They all follow the paradigm of first selecting channels or layers to prune based on a designed metric, and then performing RFT (Chavan et al.
+2024).
+However, they neglect that important information can also exist in the pruned parts (Dettmers et al.
+2022; Xiao et al.
+2023; Yin et al.
+2024), and directly removing them leads to an irreversible decline in model performance.
+Thus, the pruned models often require extensive data for RFT to recover performance (Ma, Fang, and Wang 2023).
+Additionally, high pruning ratios often cause performance collapse, limiting their effectiveness in acceleration.
+To tackle these limitations, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then prunes those parameters, finally performs RFT.
 To the best of our knowledge, we are the first to propose this paradigm.
@@
 Based on this paradigm, we introduce DReSS, a Data-driven Regularized Structured Streamlining method for LLMs.
-The comparison of the proposed method with existing channel-wise pruning approaches is shown in Figure 1.
-Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, finally (4) do RFT.
+The comparison of the DReSS with existing channel-wise pruning approaches is shown in Figure 1.
+Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrices, (3) prune the regularized channels, finally (4) do RFT.
 First, we randomly select a small subset of data from widely used benchmark datasets.
-Since the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced.
-Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization to these channels using the selected data.
-This redistributes important information to the unpruned components, thereby significantly mitigating performance degradation caused by direct pruning.
-As a result, the model can maintain robust performance even at high pruning ratios, facilitating a substantial speedup in inference.
-Third, we prune the parameters that have been regularized.
+As the data size is small, the overhead of regularization and RFT remains minimal.
+Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization ($\ell_{1}$-norm or $\ell_{2}$-norm) to these channels using the selected data.
+This redistributes important information to the unpruned parts, thereby significantly mitigating performance degradation caused by direct pruning.
+As a result, the model retains strong performance at high pruning ratios, enabling significant speedup.
+Third, we prune the regularized channels.
 Finally, we perform RFT on the pruned model using a small subset of samples selected in the first step.
-Extensive experiments show that DReSS significantly outperforms existing pruning methods in perplexity and accuracy across various pruning ratios, even without RFT, while also greatly enhancing inference speed.
-The main contributions of our work are summarized as follows:
-  • Propose A New Paradigm: By sequentially applying regularization, pruning, and RFT, DReSS minimizes information loss caused by direct parameter removal, thereby improving the model’s language modeling capabilities and reducing the computational cost of RFT.
-  • High Performance: DReSS surpasses existing structured pruning methods in both generation and zero-shot tasks.
-The pruned model achieves a substantial increase in throughput and a significant reduction in latency compared to the dense model.
-  • Low Overhead: By leveraging only a small amount of data for regularization and optional RFT, DReSS incurs minimal overhead.
-Figure 1: A comparison between existing channel-wise structured pruning methods and DReSS.
-The deeper blue square represents greater performance impact; the red exclamation mark indicates the important information.
-The taller orange-yellow cylinder represents larger data volume.
-L($\cdot$) represents the regularization, and S represents pruning.
-Above: Existing methods first select channels based on importance, followed by pruning and RFT on the pruned model.
-Below: DReSS first regularizes channels to transfer important information, prunes the channels, then do RFT with a little amount of data.
+Extensive experiments show that DReSS significantly outperforms existing powerful pruning methods in perplexity and accuracy across various pruning ratios on different models and is able to provide considerable acceleration.
+The main contributions are summarized as follows:
+  • Propose A New Paradigm: By sequentially applying regularization ($\ell_{1}$-norm or $\ell_{2}$-norm), pruning, and RFT, DReSS minimizes information loss caused by direct parameter removal, thereby improving the model’s language modeling capabilities.
+  • High Performance: DReSS surpasses competitive structured pruning methods in perplexity and accuracy, achieving notable improvements in throughput and reduced latency compared to the dense models.
+  • Low Overhead: By using only a small amount of data for regularization and optional RFT, DReSS incurs minimal overhead.
+Figure 1: A comparison between previous channel-wise pruning methods and DReSS.
+Deeper blue square represents greater performance impact, the taller cylinder represents larger data volume.
+Above: Previous methods first select channels based on importance, followed by pruning and then do RFT.
+Below: DReSS first regularizes channels to transfer important information, prunes the channels, then do RFT with a small amount of data.
 2 Related Works
 2.1 Pruning Methods
-The elimination of redundant weights, referred to as pruning, has been an effective approach for reducing the complexity of deep neural networks over the past several decades (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015; Wen et al., 2016; Louizos et al., 2017).
-Pruning methods can be broadly categorized into two types: unstructured pruning (Kurtic et al., 2022; Zhang et al., 2024; Xu et al., 2024) and structured pruning (Xia et al., 2023b; Gao et al., 2024b; Yang et al., 2024).
-Unstructured pruning methods remove individual weights based on their importance.
-For instance, Magnitude (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix.
-The main advantage is flexibility, enabling the removal of unimportant weights while retaining critical ones, allowing high sparsity in LLMs.
-However, its limitations include the requirement for specialized hardware to accelerate inference (Xia et al., 2023a), the necessity for high sparsity to achieve substantial acceleration (Wang, 2020), and the inability to perform RFT on downstream tasks.
-Structured pruning methods, including channel-wise pruning (An et al., 2024) and layer-wise pruning (Men et al., 2024), can achieve significant inference speedup with a low sparsity.
-Channel-wise pruning methods typically design a metric to evaluate the importance of channels in the parameter matrix of LLMs and then prune the less important channels.
-For instance, SliceGPT (Ashkboos et al., 2024) adopts Principal Component Analysis (PCA), using the eigenvalues of the covariance matrix of activation values as a metric to preserve channels corresponding to larger eigenvalues.
-LLM Surgeon (van der Ouderaa et al., 2023) periodically updates model weights and structures, pruning more aggressively in the initial layers and less so in the intermediate layers.
-Layer-wise pruning methods, such as SLEB (Song et al., 2024), iteratively prune entire transformer layers by assessing their importance.
-However, these methods have a key limitation: even less important channels or layers may contain valuable information, and pruning them directly often leads to significant performance degradation.
-To address this, we investigate the possibility of ‘shifting’ important information from the pruned parts to the retained parts, which could substantially improve model performance.
+Pruning redundant weights has been an effective way to reduce deep neural network complexity for decades.
+(LeCun, Denker, and Solla 1989; Hassibi, Stork, and Wolff 1993; Han et al.
+2015).
+Pruning methods can be broadly categorized into two types: unstructured pruning (Kurtic et al.
+2022; Zhang et al.
+2024; Xu et al.
+2024) and structured pruning (Xia et al.
+2024; Gao et al.
+2024b).
+Unstructured pruning methods remove individual weights based on designed metrics.
+Magnitude (Han et al.
+2015) removes smaller weights, Wanda (Sun et al.
+2024) considers both weights and activations, and SparseGPT (Frantar and Alistarh 2023) uses Hessian matrix.
+They require specialized hardware to accelerate (Xia et al.
+2023) and a high sparsity to achieve substantial acceleration (Wang 2020).
+Structured pruning methods, including channel-wise pruning (An et al.
+2024) and layer-wise pruning (Men et al.
+2024).
+Channel-wise pruning methods remove less important channels of the parameter matrices.
+For instance, SliceGPT (Ashkboos et al.
+2024) prune channels with smaller eigenvalues.
+LLM Surgeon (van der Ouderaa et al.
+2024) periodically updates model weights and structures, pruning more aggressively in the initial layers.
+Layer-wise pruning methods, such as SLEB (Song et al.
+2024), iteratively prune transformer layers based on their importance.
+These methods have a key limitation: even less important channels or layers may contain valuable information, and pruning them directly often leads to information loss.
+To address this, we investigate the way of ‘shifting’ important information from the pruned parts to the remaining parts, which could reduce information loss caused by pruning.
 2.2 Regularization
-Regularization is widely employed in machine learning (Hoerl & Kennard, 1970; Poggio et al., 1987; Balestriero et al., 2022), such as feature selection (Tibshirani, 1996) and preventing model overfitting (Santos & Papa, 2022).
-The $\ell_{1}$-norm typically induces sparsity by driving certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd & Vandenberghe, 2004).
-Both of them can significantly alter the distribution pattern of the data (Han et al., 2015; Tao et al., 2023).
-Motivated by this observation, we propose a data-driven regularization approach.
-By applying a small amount of data, the regularization transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
+Regularization is widely used in machine learning, such as feature selection (Tibshirani 1996) and preventing model overfitting (Santos and Papa 2022).
+The $\ell_{1}$-norm drives certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd and Vandenberghe 2004).
+Both of them can significantly alter the distribution pattern of the data (Han et al.
+2015; Tao et al.
+2023).
+Motivated by this, we propose a new pruning paradigm, in which the regularization process can transfer important information from the pruned parameter space to the remaining parts of the model, thus reducing the information loss caused by parameter removal.
 3 Methodology
-In this section, we provide a detailed introduction to DReSS. The overall DReSS algorithm is summarized in Algorithm 1, our method consists of four main steps:
+In this section, we provide a detailed introduction to DReSS. The overall DReSS algorithm is summarized in Figure 1 and Algorithm 1, our method consists of four main steps:
   • Data Selection: Select a small amount of data for pre-pruning regularization and post-pruning RFT.
-  • Regularization: Applying regularization to the selected channels of weight matrices during forward and backward propagations.
-  • Pruing: Prune the regularized portions.
-  • RFT: Perform RFT on the pruned model using the subset of data selected in the first step.
+  • Regularization: Apply regularization to the selected channels of weight matrices to transfer important information from the pruned parts to the remaining parts, thereby reducing information loss caused by pruning and mitigating the resulting performance degradation.
+  • Pruing: Remove the channels that were regularized in the previous step.
+  • RFT: Perform RFT on the pruned model to compensate for the performance loss caused by pruning.
+3.1 Data Selection
+For fairness, the data is selected from widely used datasets.
+For instance, approximately 1,000 samples are randomly selected from the WikiText-2 (Merity et al.
+2016) training set for pre-pruning regularization and post-pruning RFT.
+3.2 Regularization
+For clarity, we first define the notations.
+$p$ is the pruning ratio, indicating $p\%$ of model parameters will be pruned.
+The model dimension is $d$, and the number of layers is $l$.
+Additionally, $\mathbf{W_{emb}}\in\mathbb{R}^{b\times n\times d}$ is the data feature map, and $\mathbf{W_{pos}}\in\mathbb{R}^{b\times n\times d}$ is the positional feature map, where $b$ is the batch size and $n$ is the number of tokens.
+The model weights are represented by $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$.
+Specifically, in the $i$th layer attention block, $\mathbf{W_{Q}^{i}}\in\mathbb{R}^{d\times d_{1}}$ is query matrix, $\mathbf{W_{K}^{i}}\in\mathbb{R}^{d\times d_{1}}$ is key matrix, $\mathbf{W_{V}^{i}}\in\mathbb{R}^{d\times d_{1}}$ is value matrix, and $\mathbf{W_{o}^{i}}\in\mathbb{R}^{d_{1}\times d}$ is output matrix.
+In the $i$th layer FFN block, $\mathbf{W_{up}^{i}}\in\mathbb{R}^{d\times d_{2}}$ is up projection matrix, and $\mathbf{W_{down}^{i}}\in\mathbb{R}^{d_{2}\times d}$ is down projection matrix.
+In the $i$th LayerNorm, $\mathbf{\alpha^{i}}\in\mathbb{R}^{d}$ is scaling vector and $\mathbf{\beta^{i}}\in\mathbb{R}^{d}$ is offset vector.
+$\mathbf{W_{lm}}\in\mathbb{R}^{d\times b\times n}$ denotes the language modeling matrix.
+$\mathbf{I}\in\mathbb{R}^{d\times d}$ is the identity matrix, let $\mathbf{R}\in\mathbb{R}^{d\times d}$ is a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
+$\mathbf{R}$ is identical across all layers.
 Algorithm 1 DReSS algorithm.
-  Input: selected data $X$, number of layers in LLMs $l$, initial model $W$, number of regularization iterations $n$, norm type: flag.
-  if flag==1 then
-     use Proposition 3.2 to transform problem
+  Input: selected data $\mathbf{X}$, number of layers in LLMs $l$, initial model $\mathbf{W}$, norm type: flag.
+  if flag==0 then
+     continue// $\ell_{2}$-norm
+  else if flag==1 then
+     use Proposition 1 to transform problem // $\ell_{1}$-norm
   end if
-  $X_{1},X_{2}\leftarrow divide(X)$
-  for $i=1$ to $n$ do
+  $\mathbf{X_{1}},\mathbf{X_{2}}\leftarrow divide(\mathbf{X})$
+  ~$\tilde{\mathcal{L}}_{\text{1}}\leftarrow regularize(\mathbf{W_{emb}},\mathbf{W_{pos}},\mathbf{X_{1}})$
+  ~$\tilde{\mathcal{L}}_{\text{att}}\leftarrow 0$, ~$\tilde{\mathcal{L}}_{\text{FFN}}\leftarrow 0$, ~$\tilde{\mathcal{L}}_{\text{2}}\leftarrow 0$
+  for $i=1$ to $l$ do
+     ~$\tilde{\mathcal{L}}_{\text{att}}\leftarrow\tilde{\mathcal{L}}_{\text{att}}+regularize(Attention,\mathbf{X_{1}})$
+     ~$\tilde{\mathcal{L}}_{\text{FFN}}\leftarrow\tilde{\mathcal{L}}_{\text{FFN}}+regularize(FFN,\mathbf{X_{1}})$
+     ~$\tilde{\mathcal{L}}_{\text{2}}\leftarrow\tilde{\mathcal{L}}_{\text{2}}+regularize(LayerNorm,\mathbf{X_{1}})$
+  end for
+  ~$\tilde{\mathcal{L}}_{\text{3}}\leftarrow regularize(\mathbf{W_{lm}},\mathbf{X_{1}})$
+  ~$\tilde{\mathcal{L}}_{\text{remain}}\leftarrow\tilde{\mathcal{L}}_{\text{1}}+\tilde{\mathcal{L}}_{\text{2}}+\tilde{\mathcal{L}}_{\text{3}}$
+  $\mathcal{L}(\mathbf{W},\mathbf{X})\leftarrow forward(\mathbf{W},\mathbf{X_{1}})$
+  $\mathcal{L}_{\text{sum}}\leftarrow\mathcal{L}(\mathbf{W},\mathbf{X})+\tilde{\mathcal{L}}_{\text{att}}+\tilde{\mathcal{L}}_{\text{FFN}}+\tilde{\mathcal{L}}_{\text{remain}}$
+  update $\mathbf{W}$ using backpropagation algorithm
+  $\mathbf{W}\leftarrow Prune(\mathbf{W},\mathbf{R})$
+  $\mathbf{W}\leftarrow Optional\_RFT(\mathbf{W},\mathbf{X_{2}})$
+Figure 2: The regularization and pruning process of DReSS in a transformer layer.
+Proposition 1 (Dependency between attention and FFN block.
+Proof in Appendix A).
+If regularization is applied to certain columns of $\mathbf{W_{down}^{i-1}}$ in the $(i-1)$th layer FFN block, then regularization should also be applied to the corresponding rows of $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$ in the $i$th layer attention block.
+The same dependency applies to $\mathbf{W_{o}^{i}}$ and $\mathbf{W_{up}^{i-1}}$.
+Figure 2 illustrates the process of applying regularization in a transformer layer.
+Following Proposition 1, in the attention block, regularization is applied to the rows of $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$ corresponding to the one entries in the index matrix $\mathbf{R}$ and the columns of $\mathbf{W_{o}^{i}}$ associated with these entries.
+In the FFN block, regularization is applied on the rows of $\mathbf{W_{up}^{i}}$ and the columns of $\mathbf{W_{down}^{i}}$ that correspond to the one entries in $\mathbf{R}$.
+We also apply regularization to the remaining parts in LLMs.
+Specifically, for $\mathbf{W_{emb}}$, regularization is enforced on the columns corresponding to the one indices in $\mathbf{R}$, and the same for $\mathbf{W_{pos}}$.
+For each $\mathbf{\alpha^{i}}$ and $\mathbf{\beta^{i}}$ within every transformer layer, regularization is applied to the elements corresponding to the one indices in $\mathbf{R}$.
+Additionally, for the final $\mathbf{W_{lm}}$, due to its dependency on $\mathbf{W_{down}^{l}}$ in the FFN block of the last transformer layer, regularization is applied to the corresponding rows.
+Therefore, the overall loss function defined in Equation 1 consists of the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$ and the regularization loss.
+$\mathbf{X}$ is the input data.
+Regularization loss has three parts: Attention loss: $\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}$, FFN loss: $\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}$, remain loss: ~$\tilde{\mathcal{L}}_{\text{remain}}$.
+$\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(\mathbf{W},\mathbf{X})+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}\end{split}$ (1)
+~$\tilde{\mathcal{L}}_{\text{att}}^{i}$ denotes the $i$th layer attention block regularization loss:
+~$\begin{split}\tilde{\mathcal{L}}_{\text{att}}^{i}=\lambda(\|\mathbf{RW_{Q}^{i}}\|+\|\mathbf{RW_{K}^{i}}\|+\|\mathbf{RW_{V}^{i}}\|+\|\mathbf{W_{o}^{i}R}\|)\end{split}$ (2)
+~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ is the $i$th layer FFN block regularization loss:
+~$\begin{split}\tilde{\mathcal{L}}_{\text{FFN}}^{i}=\lambda\|\mathbf{RW_{up}^{i}}\|+\lambda\|\mathbf{W_{down}^{i}R}\|\end{split}$ (3)
+~$\tilde{\mathcal{L}}_{\text{remain}}$ denotes the regularization loss of the remaining parts of the LLM:
+~$\begin{split}\tilde{\mathcal{L}}_{\text{remain}}=\lambda(\|\mathbf{W_{emb}R}\|+\|\mathbf{W_{pos}R}\|+\\ \sum_{i=1}^{l}\|\mathbf{R\alpha^{i}}\|+\sum_{i=1}^{l}\|\mathbf{R\beta^{i}}\|+\|\mathbf{RW_{lm}}\|)\end{split}$ (4)
 […]
-     ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$00
-     for $j=1$ to $l$ do
-        ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{att}}$+$regularize(W_{Q}^{j},W_{K}^{j},W_{V}^{j},W_{o}^{j},X_{1})$
-        ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{FFN}}$+$regularize(W_{up}^{j},W_{down}^{j},X_{1})$
-        ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{2}}$+$regularize(w_{N}^{j},X_{1})$
-     end for
-[…]
-     ~$\tilde{\mathcal{L}}_{\text{remain}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{1}}$+~$\tilde{\mathcal{L}}_{\text{2}}$+~$\tilde{\mathcal{L}}_{\text{3}}$
-     $\mathcal{L}(W,X)\leftarrow forward(W,X_{1})$
-     $\mathcal{L}_{\text{sum}}$$\leftarrow$$\mathcal{L}(W,X)$+~$\tilde{\mathcal{L}}_{\text{att}}$+~$\tilde{\mathcal{L}}_{\text{FFN}}$+~$\tilde{\mathcal{L}}_{\text{remain}}$
-     update $W$ using backpropagation algorithm
-  end for
-  $W\leftarrow Prune(W,R)$
-  $W\leftarrow Optional\_RFT(W,X_{2})$
-3.1 Data Selection
-The required data is selected from widely used calibration datasets, including Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020).
-For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for pre-pruning regularization and post-pruning RFT.
-In Section 4.5, we analyze the impact of using different datasets.
-3.2 Regularization
-To provide a clearer understanding of this section, we first introduce the relevant notations.
-Let $p$ denote the pruning ratio, which indicates that $p\%$ of the model’s parameters need to be pruned.
-The model dimension of the LLMs is represented by $d$, and the number of layers is denoted by $l$.
-Additionally, let $W_{emb}\in\mathbb{R}^{b\times n\times d}$ represents the data feature map, and $W_{pos}\in\mathbb{R}^{b\times n\times d}$ denotes the positional feature map, where $b$ is the batch size and $n$ is the number of tokens.
-The model weights are represented by $W\in\mathbb{R}^{d_{1}\times d_{2}}$.
-Specifically, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{K}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}^{i}\in\mathbb{R}^{d\times d_{1}}$, and $W_{o}^{i}\in\mathbb{R}^{d_{1}\times d}$ represent the parameter matrices in the $i$th layer Attention block.
-Similarly, $W_{up}^{i}\in\mathbb{R}^{d\times d_{2}}$ and $W_{down}^{i}\in\mathbb{R}^{d_{2}\times d}$ represent the parameter matrices in the $i$th layer FFN block.
-Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$th layer, and $W_{lm}\in\mathbb{R}^{d\times b\times n}$ represents the language modeling matrix.
-Finally, $I\in\mathbb{R}^{d\times d}$ represents the identity matrix, let $R\in\mathbb{R}^{d\times d}$ represent a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
-Figure 2: The pruning strategy of DReSS in a transformer layer.
-Proposition 3.1.
-(Dependency between the Attention Block and FFN Block.
-Proof in Appendix A.1).
-If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block.
-The same dependency applies to $W_{o}^{i}$ and $W_{up}^{i-1}$.
-We illustrate the process of applying regularization within a transformer layer (Vaswani, 2017), as shown in Figure 2.
-Following Proposition 3.1, in the Attention block, regularization is applied to the rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ corresponding to the one entries in the index matrix $R$, as well as to the columns of $W_{o}^{i}$ associated with these entries.
-Similarly, in the FFN block, regularization is enforced on the rows of $W_{up}^{i}$ and the columns of $W_{down}^{i}$ that correspond to the one entries in $R$.
-We also apply regularization to the remaining parts in LLMs.
-Specifically, for $W_{emb}$, regularization is enforced on the columns corresponding to the one indices in $R$, and the same procedure is applied to $W_{pos}$.
-For each $w_{N}^{i}$ within every transformer layer, regularization is applied to the elements corresponding to the one indices in $R$.
-Additionally, for the final $W_{lm}$, due to its dependency on $W_{down}^{l}$ in the FFN block of the last transformer layer, regularization is applied to the corresponding rows.
-Therefore, the overall loss function consists of the language modeling loss and the loss induced by regularization, which is defined below:
-$\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(W,X)+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}\end{split}$ (1)
-where $\mathcal{L}(W,X)$ denotes the language modeling loss and $X$ is the input data, and the remaining term represents the regularization loss.
-~$\tilde{\mathcal{L}}_{i}^{\text{att}}$ denotes the $i$th layer attention block regularization loss, which is defined as:
-~$\begin{split}\tilde{\mathcal{L}}_{i}^{\text{att}}=\lambda(\|RW_{Q}^{i}\|+\|RW_{K}^{i}\|+\|RW_{V}^{i}\|+\|W_{o}^{i}R\|)\end{split}$ (2)
-~$\tilde{\mathcal{L}}_{i}^{\text{FFN}}$ denotes the $i$th layer FFN block regularization loss and the definition is as below:
-~$\begin{split}\tilde{\mathcal{L}}_{\text{FFN}}^{i}=\lambda\|RW_{up}^{i}\|+\lambda\|W_{down}^{i}R\|\end{split}$ (3)
-~$\tilde{\mathcal{L}}_{\text{remain}}$ denotes the remaining parts of the LLMs’ regularization loss as follows:
-~$\begin{split}\tilde{\mathcal{L}}_{\text{remain}}=\lambda(\|W_{emb}R\|+\|W_{pos}R\|+\\ \sum_{i=1}^{l}\|Rw_{N}^{i}\|+\|RW_{lm}\|)\end{split}$ (4)
-[…]
-Proposition 3.2.
-(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation.
-Proof in Appendix A.2).
-The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
+Proposition 2 (If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation.
+Proof in Appendix B).
+The following two problems are equivalent, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
 $\min\quad||x||_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (5)
@@
 We formalize the minimization of the overall loss as an optimization problem.
-When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation (BP) algorithm (Rumelhart et al., 1986).
-In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example.
-According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by BP algorithm.
-$\min_{W}\quad\mathcal{L}(W,X)+\lambda||RW_{up}^{i}||_{1}+\lambda||W_{down}^{i}R||_{1}$ (6)
-$\min_{W,Y_{1},Y_{2}}$ $\quad\mathcal{L}(W,X)+\mathbf{1}^{T}Y_{1}\mathbf{1}+\mathbf{1}^{T}Y_{2}\mathbf{1}$ (7)
-s.t. $\quad-Y_{1}\leq RW_{up}^{i}\leq Y_{1},$
-$\quad-Y_{2}\leq W_{down}^{i}R\leq Y_{2},$
-$\quad Y_{1}\geq 0,\quad Y_{2}\geq 0.$
-During the forward and backward propagation processes, due to the application of regularization on certain specific channels in LLMs, the values within these rows or columns are significantly reduced.
-Intuitively, this reduction could potentially transfer important information and decrease the impact of these parameters on the model’s performance.
+When $\ell_{2}$-norm is used, the problem can be directly solved using backpropagation (BP) algorithm (Rumelhart, Hinton, and Williams 1986).
+In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example, the complete objective function transformation is in Appendix C.
+According to Proposition 2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by BP algorithm.
+During forward and backward propagation, regularization on certain channels in LLMs significantly reduces the values in those rows or columns.
+Intuitively, this value reduction could decrease the impact of these parameters on the model’s performance thus potentially transfer important information to the unregularized parts.
+$\min_{\mathbf{W}}\hskip 5.0pt\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda||\mathbf{RW_{up}^{i}}||_{1}+\lambda||\mathbf{W_{down}^{i}R}||_{1}$ (6)
+$\min_{\mathbf{W,Y_{1},Y_{2}}}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda(\mathbf{1}^{T}\mathbf{Y_{1}}\mathbf{1}+\mathbf{1}^{T}\mathbf{Y_{2}}\mathbf{1})$ (7)
+s.t. $\quad-\mathbf{Y_{1}}\leq\mathbf{RW_{up}^{i}}\leq\mathbf{Y_{1}},$
+$\quad-\mathbf{Y_{2}}\leq\mathbf{W_{down}^{i}R}\leq\mathbf{Y_{2}},$
+$\quad\mathbf{Y_{1}}\geq 0,\quad\mathbf{Y_{2}}\geq 0.$
 3.3 Pruning
-After completing the regularization, we prune the rows and columns directly using the pseudo-index selection matrix $R$.
-For clarity, we define $S=I-R$ and the pruning for the $i$th layer attention block is then performed as:
-${W_{Q}^{i}}^{\prime}$ $=SW_{Q}^{i},\quad{W_{K}^{i}}^{\prime}=SW_{K}^{i}$ (8)
-${W_{V}^{i}}^{\prime}$ $=SW_{V}^{i},\quad{W_{o}^{i}}^{\prime}=W_{o}^{i}S$
-While for the FFN block, it follows:
-${W_{up}^{i}}^{\prime}=SW_{up}^{i},\quad{W_{down}^{i}}^{\prime}=W_{down}^{i}S$ (9)
+According to Figure 2, we prune the regularized rows and columns using the pseudo-index selection matrix $\mathbf{R}$.
+For clarity, we define $\mathbf{S=I-R}$.
+The pruning of the $i$th layer attention block and FFN block are performed as:
+$\mathbf{{W_{Q}^{i}}^{\prime}}$ $\mathbf{=SW_{Q}^{i},}\quad$ $\mathbf{{W_{K}^{i}}^{\prime}}$ $\mathbf{=SW_{K}^{i},}$ (8)
+$\mathbf{{W_{V}^{i}}^{\prime}}$ $\mathbf{=SW_{V}^{i},}\quad$ $\mathbf{{W_{o}^{i}}^{\prime}}$ $\mathbf{=W_{o}^{i}S,}$
+$\mathbf{{W_{up}^{i}}^{\prime}}$ $\mathbf{=SW_{up}^{i},}\quad$ $\mathbf{{W_{down}^{i}}^{\prime}}$ $\mathbf{=W_{down}^{i}S}$
 The remaining parts are pruned as below:
-${W_{emb}}^{\prime}$ $=W_{emb}S,\quad{W_{pos}}^{\prime}=W_{pos}S$ (10)
-${w_{N}^{i}}^{\prime}$ $=S{w_{N}^{i}},\quad{W_{lm}}^{\prime}=SW_{lm}$
-The entire pruning process is straightforward and simple.
+$\mathbf{{W_{emb}}^{\prime}=W_{emb}S,\quad{W_{pos}}^{\prime}=W_{pos}S,}$ (9)
+$\mathbf{{\alpha^{i}}^{\prime}=S{\alpha^{i}},\quad{\beta^{i}}^{\prime}=S{\beta^{i}},\quad{W_{lm}}^{\prime}=SW_{lm}}$
+The pruning process is straightforward and simple.
 3.4 Optional RFT
-After pruning, we perform RFT on the model using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al., 2021).
-This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance in terms of both perplexity and zero-shot accuracy.
+After pruning, we perform RFT using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al.
+2022).
+This step is optional, as experimental results in section 4.7 show that DReSS without RFT still achieves competitive performance in perplexity and zero-shot accuracy.
+4 Experiments
+Model Method PR PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc
+Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24
+LLM Surgeon 25% 7.26 67.28 63.25 54.24 51.62 35.85 54.44
+SliceGPT 25% 7.06 69.32 65.39 52.57 53.78 31.89 54.59
+SLEB 25% 7.82 67.94 62.79 49.80 48.55 29.34 51.68
+DReSS-$\ell_{2}$ 25% 6.25 68.52 66.71 56.73 52.78 37.63 56.47
+DReSS-$\ell_{1}$ 25% 6.28 68.67 65.32 57.16 52.39 37.28 56.16
+LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00
+LLM Surgeon 25% 7.38 70.59 65.87 58.66 63.65 38.33 59.42
+SliceGPT 25% 7.49 68.15 64.13 56.22 55.39 34.74 55.73
+SLEB 25% 10.24 63.25 62.36 53.77 55.82 32.24 53.49
+DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54
+DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20
+LLaMA3-8B Dense 0% 5.76 85.56 77.94 79.27 78.84 56.49 75.62
+LLM Surgeon 25% 7.62 76.34 70.18 71.46 69.65 49.22 67.37
+SliceGPT 25% 8.14 74.37 67.81 69.56 69.83 45.53 65.42
+SLEB 25% 11.37 71.68 62.96 66.37 64.61 43.28 61.78
+DReSS-$\ell_{2}$ 25% 6.09 78.29 71.17 73.28 72.64 53.62 69.80
+DReSS-$\ell_{1}$ 25% 6.12 78.86 70.28 73.85 72.37 53.02 69.68
+OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79
+LLM Surgeon 25% 11.02 74.18 64.33 65.37 60.96 34.98 59.96
+SliceGPT 25% 10.90 73.72 64.28 63.33 60.59 34.66 59.32
+SLEB 25% 12.02 72.62 63.96 62.79 59.21 34.12 58.50
+DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47
+DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07
+LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76
+LLM Surgeon 25% 5.75 77.75 69.62 74.31 72.83 43.52 67.61
+SliceGPT 25% 6.16 69.69 68.45 63.73 63.46 39.90 61.05
+SLEB 25% 7.39 66.93 65.87 55.48 60.06 35.14 56.70
+DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84
+DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
 Table 1: Performance comparison of different pruning methods.
 ‘PPL’ refers to the perplexity on WikiText-2.
-The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy.
-The best result is highlighted in bold, and the second-best in underlined.
+The accuracy (%) is reported on five zero-shot benchmarks.
+The best result is highlighted in bold, and the second-best is underlined.
 DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
-Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 LLM Surgeon 25% 7.26 67.28 63.25 54.24 51.62 35.85 54.44 SliceGPT 25% 7.08 69.21 65.35 52.40 53.70 31.66 54.46 SLEB 25% 7.82 67.94 62.79 49.80 48.55 29.34 51.68 DReSS-$\ell_{2}$ 25% 6.25 68.52 66.71 56.73 52.78 37.63 56.47 DReSS-$\ell_{1}$ 25% 6.28 68.67 65.32 57.16 52.39 37.28 56.16 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 LLM Surgeon 25% 12.60 68.22 58.06 46.25 51.52 27.31 50.27 SliceGPT 25% 14.56 65.29 57.22 47.85 49.79 27.99 49.63 SLEB 25% 16.73 63.28 55.37 43.09 45.62 24.21 46.31 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 LLM Surgeon 25% 11.33 72.12 61.08 62.36 54.19 30.52 56.05 SliceGPT 25% 11.90 70.35 60.62 58.15 52.78 29.52 54.28 SLEB 25% 13.80 71.28 60.96 59.26 51.54 29.83 54.57 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 LLM Surgeon 25% 11.02 74.18 64.33 65.37 60.96 34.98 59.96 SliceGPT 25% 10.94 73.67 64.25 63.28 60.52 34.64 59.27 SLEB 25% 12.02 72.62 63.96 62.79 59.21 34.12 58.50 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 LLM Surgeon 25% 7.38 70.59 65.87 58.66 63.65 38.33 59.42 SliceGPT 25% 7.55 66.87 63.38 54.16 58.46 34.56 55.48 SLEB 25% 10.24 63.25 62.36 53.77 55.82 32.24 53.49 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 LLM Surgeon 25% 5.75 77.75 69.62 74.31 72.83 43.52 67.61 SliceGPT 25% 6.30 68.55 67.48 58.10 62.50 37.88 58.90 SLEB 25% 7.39 66.93 65.87 55.48 60.06 35.14 56.70 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
-Figure 3: The ratio of the sum of absolute parameter values after regularization to that before regularization for LLaMA2-7B.
-For example, regularization is applied to the last 25% of rows in $W_{Q}$, and the absolute sum of these rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Preserved Ratios.
-This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the regularized columns is compared to the original sum.
-4 Experiments
-This section introduces experimental setup (4.1) and analyzes the effectiveness of DReSS from the following aspects: performance comparison (4.2), acceleration (4.3), robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (4.7).
+The results below are all obtained after RFT.
 4.1 Experimental Setup
-Implementation: DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019) , utilizing the Hugging Face Transformers library (Wolf, 2019).
+Implementation: All methods are implemented in PyTorch (Paszke et al.
+2019), using the Hugging Face Transformers library (Wolf 2019).
 All experiments are conducted on 80GB NVIDIA A100 GPUs.
-For fairness, we use llm-eval-harness (Gao et al., 2024a) to evaluate the pruned models.
-More details are shown in Appendix B.
-Datasets: To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks.
-For generation task, following prior work(Ashkboos et al., 2024), we evaluate the model’s perplexity on WikiText-2.
-For zero-shot task evaluation, the benchmarks consists of PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018).
-To demonstrate that DReSS is not specifically dependent on the regularization data, we use a randomly selected subset of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration before pruning and for RFT after pruning.
-Models: The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
-Baselines: We evaluate DReSS against competitive structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024).
-Among them, LLM Surgeon and SliceGPT are channel-wise, SLEB is layer-wise.
-To ensure a fair comparison, in the subsequent experiments, we ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB before pruning, and that the data used for RFT in each method is also the same.
-We also compared with powerful unstructured pruning methods like Mangniude (Han et al., 2015), Wanda (Sun et al., 2023), and SparseGPT (Frantar & Alistarh, 2023) and the results are in Appendix C.
-Evaluation Metrics: The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024).
-The acceleration effects are represented by throughput and latency (Song et al., 2024).
+For fairness, we use llm-eval-harness (Gao et al.
+2024a) to evaluate.
+Datasets: For generation task, we evaluate the model’s perplexity on WikiText-2 (Ashkboos et al.
+2024).
+For zero-shot task, the benchmarks are PIQA (Bisk et al.
+2020), WinoGrande (Sakaguchi et al.
+2021), HellaSwag (Zellers et al.
+2019), ARC-e and ARC-c (Clark et al.
+2018).
+In Appendix H we use data from Alpaca (Taori et al.
+2023), WikiText-2 (Merity et al.
+2016), PTB (Marcus, Santorini, and Marcinkiewicz 1993), and C4 (Raffel et al.
+2020).
+Models: We use models commonly adopted in the pruning domain including the LLaMA models (LLaMA2-7B, LLaMA3-8B, LLaMA2-13B) (Touvron et al.
+2023; Grattafiori et al.
+2024), OPT model (OPT-13B) (Zhang et al.
+2022), and Phi-2 (Javaheripi et al.
+2023).
+Baselines: We evaluate DReSS against competitive structured pruning methods: LLM Surgeon (van der Ouderaa et al. 2024), SliceGPT (Ashkboos et al. 2024), and SLEB (Song et al. 2024).
+Evaluation Metrics: The performance on generation task is measured by perplexity(Yao et al.
+2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al.
+2024).
+The acceleration effects are represented by throughput and latency (Song et al.
+2024).
+Model Method PR PPL($\downarrow$) Tokens/s($\uparrow$) TI($\uparrow$) Latency($\downarrow$) Speedup($\uparrow$)
+OPT-13B Dense 0% 10.12 1029 1.00× 386.5 1.00×
+DReSS 25% 10.38 1194 1.16× 319.42 1.21×
+DReSS 50% 13.85 1389 1.35× 274.11 1.41×
+LLaMA2-13B Dense 0% 4.88 1066 1.00× 396.9 1.00×
+DReSS 25% 5.12 1215 1.14× 330.8 1.20×
+DReSS 50% 7.59 1407 1.32× 285.5 1.39×
+Table 2: Comparison of throughput and latency under different ratios on OPT-13B and LLaMA2-13B.
+‘PPL’ refers to the perplexity on Wikitext2, ‘PR’ represents ‘pruning ratio’, ‘TI’ represents ‘throughput increase’.
 4.2 Performance Comparison
-To ensure fairness, we randomly selected 1,000 samples from WikiText-2 for each method with a sequence length of 2048.
-The ratio of calibartion data and the RFT data was set to 3:1, and the pruning ratio was 25%.
-As shown in Table 1, DReSS demonstrates superior performance across both generation and zero-shot tasks achieving lowest perplexity and highest average accuracy across all models.
-For instance, on LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon.
-Notably, on OPT-13B, compared to the dense model, DReSS experiences only a 1% drop in average accuracy.
-The performance difference between the $\ell_{2}$-norm and $\ell_{1}$-norm is tiny.
-In Figure 3, we present the ratio of the absolute values of model parameters and activation values after and before regularization of the selected portions.
-The results show that, in most intermediate layers, the parameters and activation values after regularization are reduced to 30% of their original values, which indicates that the amount of information contained in this part of the parameter space has decreased.
-We also present the changes occurring in the unregularized part in Appendix D.
-The results show that this part has increased compared to before, indicating that it now accommodates more important information.
-Overall, regularization transfers important information hidden in the pruned portions to the remaining parts.
+For fairness, we randomly selected 1,000 samples from WikiText-2 training set with sequence length of 2048 for each method.
+We ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB, and the data used for RFT in each method is also the same.
+The ratio of calibartion and RFT data was set to 3:1, the reason is discussed in Section 4.7.
+All methods are implemented under this setup.
+The pruning ratio was 25%.
+The last 25% of the values in the diagonal matrix $\mathbf{R}$ are set to 1, which means pruning the last 25% of the rows or columns of the parameter matrix and we discussed the choice of $\mathbf{R}$ in Section 4.4.
+As shown in Table 1, DReSS achieves superior performance in both generation and zero-shot tasks.
+On LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon.
+On OPT-13B, compared to the dense model, DReSS only drops 1% in average accuracy.
+The performance difference between $\ell_{2}$-norm and $\ell_{1}$-norm is tiny.
+In the following, DReSS refers to the $\ell_{2}$-norm version.
 4.3 Acceleration Effectiveness
@@
 We separately analyze the speedup achieved in each stage.
-Table 2 presents the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated using a single NVIDIA A100 GPU.
-Following the methodology of previous work (Song et al., 2024), the token generation test scenario involves generating sentences of 128 tokens with a batch size of 64, whereas for prompt processing, latency is measured by processing an input sequence of 2048 tokens.
+Table 2 presents the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated using a single 80GB NVIDIA A100 GPU.
+Following the methodology of previous work (Song et al.
+2024), the token generation test involves generating sentences of 128 tokens with a batch size of 64, whereas for prompt processing, latency is measured by processing an input sequence of 2048 tokens.
 At a pruning ratio of 50% on OPT-13B, DReSS delivers a 35% improvement in throughput and a 30% reduction in latency compared to the dense model.
-These results highlight the superior efficiency of DReSS in accelerating model inference.
-Table 2: Comparison of throughput and latency under different ratios on OPT-13B and LLaMA2-13B.
-‘PPL’ refers to the perplexity on Wikitext2, ‘PR’ represents ‘pruning ratio’, ‘TI’ represents ‘throughput increase’.
-Model Method PR PPL($\downarrow$) Tokens/s TI Latency(ms) Speedup OPT-13B Dense 0% 10.12 1029 1.00× 386.5 1.00× DReSS 25% 10.38 1194 1.16× 319.42 1.21× DReSS 50% 12.15 1389 1.35× 274.11 1.41× LLaMA2-13B Dense 0% 4.88 1066 1.00× 396.9 1.00× DReSS 25% 5.12 1215 1.14× 330.8 1.20× DReSS 50% 7.59 1407 1.32× 285.5 1.39×
-4.4 Robustness to Varying Pruning Ratios
+These results highlight the superior efficiency of DReSS in acceleration.
+Cases (1) (2) (3)
+PPL($\downarrow$) 5.86 5.89 5.87
+Avg_Acc(%) 61.54 61.39 61.46
+Table 3: The impact of pruning different channels on LLaMA2-7B. ‘PPL’ is the perplexity on WikiText-2. Avg_Acc is on five zero-shot benchmarks.
+4.4 The Impact of Pruning Different Channels
+Matrix $\mathbf{R}$ is used to select the channels on which regularization is applied (followed by pruning).
+We choose three options of $\mathbf{R}$ under 25% sparsity: (1) remove the matrix’s last 25% rows or columns, (2) remove the first 25% rows or columns , (3) divide the matrix into 5 segments, removing each segment’s 5% of the last rows or columns.
+The results on LLaMA2-7B are shown in Table 3.
+The performance differences among (1), (2), and (3) are minimal, demonstrating that DReSS is insensitive to pruning different channels.
+4.5 Robustness to Different Pruning Ratios
 Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%.
-The perplexity of LLaMA2-7B on WikiText-2 under different methods are shown in Figure 4.
+The perplexity of LLaMA2-7B under different methods are shown in Figure 3.
 DReSS significantly outperforms other methods across various pruning ratios.
-At a pruning ratio of 20%, DReSS performs similarly to the dense model.
 When the pruning ratio is up to 60%, SLEB collapses, while DReSS maintains relatively low perplexity compared to other pruning methods.
-This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to achieve high sparsity levels and unlocking significant potential for model acceleration.
-A more detailed exploration of pruning ratios using DReSS is provided in the Appendix E.
-Figure 4: Perplexity of LLaMA-7B pruned by various approaches under different pruning ratios on WikiText-2.
-4.5 Dependency on Datasets
-Since DReSS relies on data-driven regularization, we investigate its potential dataset dependence.
-We evaluated perplexity for four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4.
-To ensure fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2.
-As shown in Figure 5, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness and scalability.
-More details are shown in Appendix F.
-Figure 5: Comparison of perplexity on different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
+This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to unlock significant potential for model acceleration.
+More detailed results are in Appendix G.
+Figure 3: Perplexity of LLaMA2-7B pruned by various approaches under different pruning ratios on WikiText-2.
+Figure 4: Perplexity of DReSS and other methods on WikiText-2 under different amount of data.
+Figure 5: The ratio of the sum of absolute parameter values after regularization to that before regularization on LLaMA2-7B.
+For example, the absolute sum of the last 25% of rows in $\mathbf{W_{Q}}$ after regularization is divided by the sum of the corresponding rows before regularization to obtain the preserved ratios.
+This is applied similarly to other matrices $\mathbf{W_{K},W_{V},W_{o},W_{gate},W_{up},W_{down}}$, and layer activations.
+Cases N&N&N R&N&N R&N&R N&N&R R&P&N R&P&R N&P&N N&P&R
+PPL($\downarrow$) 5.47 5.06 4.86 5.06 5.97 5.63 22.38 11.45
+Avg_Acc(%) 69.00 70.39 71.23 70.39 59.56 62.35 47.25 52.73
+Table 4: The model performance under different regularization settings.
+‘PPL’ is the perplexity on WikiText-2.
+Avg_Acc is on five zero-shot benchmarks.
+Pruning ration is 25% and the LLM is LLaMA2-7B.
+We didn’t perform RFT after pruning.
 4.6 Minimal Overhead
 We evaluated the perplexity of each method on LLaMA2-7B using varying amounts of data, keeping all other conditions consistent with Section 4.2 and only change the data size from 500 to 8,000.
-As shown in Table 3, when only 1,000 samples were used for regularization and RFT, DReSS outperformed the other methods that used 4,000 samples.
-This further highlights the effectiveness and efficiency of applying regularization prior to pruning, as it transfers critical information from the pruned segments in advance, enabling DReSS to deliver strong performance with minimal data overhead.
-Table 3: The impact of the amount of data on model performance.
-Data Size 500 1000 2000 4000 8000 LLM Surgeon 8.14 7.38 6.55 5.89 5.56 SliceGPT 8.63 7.55 6.82 6.07 5.74 SLEB 13.26 10.24 9.67 8.79 7.83 DReSS 6.75 5.86 5.61 5.48 5.43
+As shown in Figure 4, when only 1,000 samples were used for regularization and RFT, DReSS outperformed the other methods that used 4,000 samples.
+This further highlights the effectiveness and efficiency of applying regularization prior to pruning, as it transfers critical information in advance, enabling DReSS to get strong performance with minimal data overhead.
 4.7 Ablation Study
-Effectiveness of Regularization: As shown in Table 4, both random pruning and full parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio.
-This highlights the importance of applying regularization exclusively to the pruned components.
+As shown in Table 5, both pruning without regularization and full parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at 25% sparsity.
+This highlights the importance of applying regularization precisely to the pruned components.
 Additionally, RFT after pruning has minimal impact on model performance, indicating that the RFT process is optional.
-Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. ‘w/o’ denotes the removal of specific components.
-Model Method Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B DReSS 5.86 0.00 61.54 0.00 w FPR 12.68 +6.82 50.26 -11.28 w/o R 22.38 +16.52 47.25 -14.29 w/o RFT 5.97 + 0.11 59.56 -1.98 LLaMA2-13B DReSS 5.12 0.00 68.84 0.00 w FPR 10.53 +5.41 53.28 -15.56 w/o R 18.94 +13.82 50.17 -18.67 w/o RFT 5.39 + 0.27 67.96 -0.88
-Table 5: The perplexity on WikiText-2 under different regularization and RFT data ratio using LLaMA2-7B at a pruning ratio of 25%.
-Ratio (R/F) 4:1 3:1 2:1 1:1 1:2 1:3 PPL($\downarrow$) 5.89 5.86 6.07 6.73 7.45 8.62
-Table 6: The perplexity of different $\lambda$ on LLaMA2-7B.
-$\lambda$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ $5\times 10^{-2}$ $10^{-1}$ PPL($\downarrow$) 7.47 6.04 5.86 5.93 6.12 7.26 10.55
-Data Ratio of Regualrization and RFT: With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5.
-It can be observed that when the ratio is 3:1, the model achieves the lowest perplexity.
-When the ratio is 4:1, the increase in perplexity is less pronounced, but when the ratio is 1:1, the perplexity increases significantly by 15%.
-These results underscore the importance of the regularization process relative to the RFT applied after pruning.
+Data Ratio of Regualrization and RFT: With other settings consistent with Section 4.2, we only vary the data ratio $[4:1,3:1,2:1,1:1,1:2,1:3]$.
+The lowest perplexity is achieved under 3:1 ratio.
+We speculate that pre-pruning regularization may be more critical than post-pruning RFT.
+Model Setting PPL($\downarrow$) AVG_ACC
+LLaMA2-7B DReSS 5.86 61.54
+w FPR 12.68 50.26
+w/o R 22.38 47.25
+w/o RFT 5.97 59.56
+LLaMA2-13B DReSS 5.12 68.84
+w FPR 10.53 53.28
+w/o R 18.94 50.17
+w/o RFT 5.39 67.96
+Table 5: Ablation results on LLaMA2-7B, LLaMA2-13B. ‘FPR’ is full parameter regularization, ‘R’ is regularization on selected channels. ‘w/o’ means removing specific parts.
 Trade-Off Between Language Modeling and Regularization: Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$.
-As shown in Table 6, when $\lambda=10^{-5}$, the regularization loss is too small to effectively shift important values, leading to a significant increase in perplexity after pruning.
-In contrast, when $\lambda=10^{-1}$, excessive regularization compromises the model’s language modeling ability, causing a notable performance decline post-pruning.
-Optimal performance is achieved at $\lambda=10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss for effective pruning.
+Optimal $\lambda$ is $10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss.
+We listed the best $\lambda$ for each model in Appendix E.
+4.8 Impact of Regularization
+In Figure 5, we present the ratio of model weights and activation values after and before regularization of the selected portions.
+We find that in most intermediate layers, the weights and activation values after regularization are reduced to 30% of their original values, which indicates that information contained in this parameter space has decreased.
+We also present the changes in the unregularized part in Appendix F.
+The results show that this part has increased compared to before, indicating that it now contains more information showing that regularization transfers important information from the pruned parts to the remaining parts.
+To evaluate regularization effectiveness, we divided into three stages: whether to regularize before pruning, whether to prune, and whether to regularize after pruning, totally 8 cases.
+For example,“N&P&R” denotes no regularization, pruning, then regularized.
+We used 25% sparsity LLaMA2-7B.“N&N&R” and “R&N&N” are the same.
+The 750 (0.75$\times$1,000) samples used for regularization are consistent with Section 4.2.
+As shown in Table 4, “R&P&N” and “N&P&R” outperform “N&P&N”, “R&P&R” is the best, showing that regularization improves performance before, after pruning, and both of them.
+“R&P&N” outperforms “N&P&R”, indicating that the paradigm of applying regularization before pruning may transfer important information to the remaining parts, thereby preserving model capacity.
 5 Conclusion
-In this paper, we introduce a novel pruning paradigm: first applying regularization to the pruned components, followed by pruning and then RFT.
-Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during regularization, effectively mitigating the irreversible performance degradation caused by information loss.
-DReSS demonstrates superior performance in both generation and zero-shot tasks, significantly outperforming existing pruning methods.
-For instance, DReSS surpasses the powerful LLM Surgeon by 20% in perplexity on LLaMA2-7B.
-On OPT-13B, with a pruning ratio of 25%, the average accuracy drops by only 1%, while achieving a 1.41× speedup compared to the dense model.
-Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing both data and computational costs while minimizing the reliance on RFT.
-Impact Statement
-Pruning plays a crucial role in model compression.
-DReSS is an efficient and low-overhead pruning method, which enables smaller models to achieve competitive performance, especially in resource-constrained scenarios and it also provides some insights for the development of LLMs.
-Appendix A Proofs of the Propositions
-A.1 Proof of Proposition 3.1
-Let us first assume that $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times k}$.
-We can express $A$ as $\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}$, where $\mathbf{a}_{i}\in\mathbb{R}^{m\times 1}$, and $B$ as $\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}$, where $\mathbf{b}_{i}\in\mathbb{R}^{k\times 1}$.
-Then, we obtain $C\in\mathbb{R}^{m\times k}$ as shown in Equation 11.
+In this paper, we propose a new pruning paradigm: apply regularization, prune, and finally RFT.
+Unlike previous paradigm that first prune and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining parts during regularization, effectively mitigating the irreversible performance degradation caused by information loss.
+DReSS demonstrates superior performance in generation and zero-shot tasks, significantly outperforming existing methods.
+For instance, DReSS surpasses the powerful LLM Surgeon by 21% in perplexity on LLaMA2-7B.
+On OPT-13B, under 25% sparsity, the average accuracy drops by only 1%, while achieving 1.21× speedup compared to the dense model.
+Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing computational costs and the reliance on RFT.
+The new paradigm may provide insights for structured pruning in LLMs.
+Appendix A Proof of the Proposition 1
+Let us first assume that $\mathbf{A}\in\mathbb{R}^{m\times n}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$.
+We can express $\mathbf{A}$ as $\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}$, where $\mathbf{a}_{i}\in\mathbb{R}^{m\times 1}$, and $\mathbf{B}$ as $\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}$, where $\mathbf{b}_{i}\in\mathbb{R}^{k\times 1}$.
+Then, we obtain $\mathbf{C}\in\mathbb{R}^{m\times k}$ as shown in Equation 10.
 It is important to note that the final summation in Equation 10 refers to the summation of $n$ rank-1 matrices, not a scalar summation.
-$\begin{split}C=AB=\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}\times\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}=\sum_{i=1}^{n}\mathbf{a}_{i}\mathbf{b}_{i}^{T}\end{split}$ (11)
-According to Equation 11, when some columns of $A$ are zero, the rows of $B$ that correspond to the outer product with the zero columns of $A$ can take any values, as the product of zero and any value is zero.
-These rows can therefore be set to zero.
-Applying the above conclusion to the parameters of a transformer layer, if certain columns of $W_{down}^{i-1}$ in the FFN layer of the $(i-1)$th transformer layer are pruned, then, since $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th transformer layer are immediately multiplied by the result of $W_{down}^{i-1}$, the corresponding rows in $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ can be set to zero.
-Additionally, if certain columns of $W_{o}^{i}$ in the $i$th transformer layer are pruned, the $W_{up}^{i}$ in the FFN block will immediately multiply with it, so the corresponding rows in $W_{up}^{i}$ should also be pruned.
-Thus, we have completed the proof of Proposition 3.1.
-A.2 Proof of Proposition 3.2
+$\begin{split}\mathbf{C}=\mathbf{AB}=\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}\times\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}=\sum_{i=1}^{n}\mathbf{a}_{i}\mathbf{b}_{i}^{T}\end{split}$ (10)
+According to Equation 10, To keep the elements of $\mathbf{a}_{i}\mathbf{b}_{i}^{T}$ small, both $\mathbf{a}_{i}$ and $\mathbf{b}_{i}^{T}$ must be small.
+So, if regularization is applied to $\mathbf{a}_{i}$ and it should also be applied to the corresponding row $\mathbf{b}_{i}^{T}$ in $\mathbf{B}$.
+Applying the above conclusion to the parameters of a transformer layer, if regularization is applied to certain columns of $\mathbf{W_{down}^{i-1}}$ in the FFN layer of the $(i-1)$th transformer layer, then, since $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$ in the $i$th transformer layer are immediately multiplied by the result of $\mathbf{W_{down}^{i-1}}$, regularization should also be applied to the corresponding rows in $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$.
+Additionally, if regularization is applied to certain columns of $\mathbf{W_{o}^{i}}$ in the $i$th transformer layer, the $\mathbf{W_{up}^{i}}$ in the FFN block will immediately multiply with it, so regularization should also be applied to the corresponding rows in $\mathbf{W_{up}^{i}}$.
+Thus, we have completed the proof of Proposition 1.
+Appendix B Proof of Proposition 2
 Step 1: Expressing $\ell_{1}$-norm Using Elements.
@@
 This transformation allows the optimization problem to be solved without directly involving the absolute value function, resulting in an equivalent constrained optimization problem that can be addressed via backpropagation.
-Thus, the proof of the Proposition 3.2 is complete.
-Appendix B Detailed Implementation
+Thus, the proof of the Proposition 2 is complete.
+Appendix C Completed Training Objective Function
+The final training objective has two parts: language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$, regularization loss.
+Regularization loss has three parts: Attention loss, FFN loss, remain loss.
+In Equation 7, for clarity, we only list one FFN layer, and principles for the others are similar.
+By using Proposition 2, the problem can be equivalently transformed into a constrained and differentiable optimization problem, which can be directly solved using the BP algorithm.
+The final objective function and constraints are as follows:
+$\mathcal{L}_{\text{sum}}$ $=\mathcal{L}(\mathbf{W,X})+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}$ (14)
+~$\tilde{\mathcal{L}}_{\text{att}}^{i}$ $=\lambda(\mathbf{\mathbf{1}^{T}Y_{3}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{4}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{5}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{6}^{i}\mathbf{1}})$
+~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ $=\lambda(\mathbf{\mathbf{1}^{T}Y_{1}^{i}\mathbf{1}+\mathbf{1}^{T}Y_{2}^{i}\mathbf{1}})$
+~$\tilde{\mathcal{L}}_{\text{remain}}$ $=\lambda\Big{(}\mathbf{1}^{T}\mathbf{Y_{7}}\mathbf{1}+\mathbf{1}^{T}\mathbf{Y_{8}}\mathbf{1}+\sum_{i=1}^{l}\mathbf{1}^{T}y_{9}^{i}\mathbf{1}+\sum_{i=1}^{l}\mathbf{1}^{T}y_{10}^{i}\mathbf{1}+\mathbf{1}^{T}\mathbf{Y_{11}}\mathbf{1}\Big{)}$
+[…]
+$-\mathbf{Y_{4}^{i}}$ $\mathbf{\leq RW_{K}^{i}\leq Y_{4}^{i},\quad-Y_{5}^{i}\leq RW_{V}^{i}\leq Y_{5}^{i},\quad-Y_{6}^{i}\leq W_{o}^{i}R\leq Y_{6}^{i}},$
+$-\mathbf{Y_{7}}$ $\mathbf{\leq W_{emb}R\leq Y_{7},\quad-Y_{8}\leq W_{pos}R\leq Y_{8}},\quad-y_{9}^{i}\mathbf{\leq R\alpha^{i}}\leq y_{9}^{i},$
+$-y_{10}^{i}$ $\mathbf{\leq R\beta^{i}}\leq y_{10}^{i},\quad-\mathbf{Y_{10}\leq RW_{lm}\leq Y_{10}}$
+$\mathbf{Y_{1}^{i}}$ $\mathbf{\geq 0,\quad Y_{2}^{i}\geq 0,\quad Y_{3}^{i}\geq 0,\quad Y_{4}^{i}\geq 0,\quad Y_{5}^{i}\geq 0,\quad Y_{6}^{i}\geq 0},$
+$\mathbf{Y_{7}}$ $\mathbf{\geq 0,\quad Y_{8}\geq 0},\quad y_{9}^{i}\geq 0,\quad y_{10}^{i}\geq 0,\quad\mathbf{Y_{11}}\geq 0$
+Appendix D Detailed Implementation
 In this part, we first introduce several hyperparameter settings, with the detailed results shown in Table 6.
-In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, OPT-6.7B, OPT-13B, LLaMA2-7B, and LLaMA2-13B.
+In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, LLaMA3-8B, OPT-13B, LLaMA2-7B, and LLaMA2-13B.
 For all RFT configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048.
-All other hyperparameters follow the default settings provided in the Hugging Face PEFT package (Mangrulkar et al., 2022).
-Due to limited computational resources, we set the batch size to 64.
-In future work, we will further explore the impact of a broader range of batch sizes on the performance of the pruning method.
-These choices ensure a balance between model adaptability and computational efficiency, leveraging parameter-efficient fine-tuning while maintaining stable optimization dynamics.
+All other hyperparameters follow the default settings provided in the Hugging Face PEFT package (Mangrulkar et al.
+2022).
+We set the batch size to 64.
+In future work, we will further explore a broader range of batch sizes.
+LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold
+32 10 2048 64 2e-5 5
 Table 6: Implementation Details
-Precision LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold Min Delta FP16 32 10 2048 64 2e-5 5 1e-4
 To ensure a fair comparison between DReSS and other methods, we maintain consistency in the data used across all approaches.
@@
 Furthermore, we ensure that the data employed during the RFT process is consistent across all methods, thereby enabling a controlled and equitable evaluation framework.
-Following previous works (Ashkboos et al., 2024; Song et al., 2024), for the comparison unstructured pruning methods like Mangniude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by DReSS for regularization.
-Appendix C Results of Comparison with Unstructured Pruning Methods
-We compare DReSS at a 25% pruning ratio with unstructured pruning methods employing 2:4 sparsity (where two out of every four consecutive parameters are set to zero) and present the experimental results in Table 8.
-For unstructured pruning methods, only calibration data is required.
-To ensure consistency, all methods utilize a total of 1,000 samples, randomly selected from WikiText-2.
-The models used in these experiments consists of the OPT and LLaMA model families.
-Compare to structured pruning methods, the primary advantage of unstructured pruning lies in its flexibility, enabling the removal of specific weights while preserving critical information in the remaining parameters of LLMs.
-Based on the results in Table 8, it can be seen that DReSS with a pruning ratio of 25% significantly outperforms existing unstructured pruning methods with a sparsity of 2:4 on generative and zero-shot tasks, demonstrating the powerful performance of DReSS.
-The reason DReSS outperforms unstructured pruning in comparisons lies in the regularization process before pruning.
-This process transfers important information hidden in the parameter space to the remaining parts of the model, preventing irreversible performance degradation caused by direct pruning and the resulting loss of information.
-Table 8: Performance comparison with unstructured pruning methods.
-‘PPL’ refers to the perplexity on WikiText-2.
-The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy.
-The best result is highlighted in bold, and the second-best in underlined.
-DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
-Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 Magnitude 2:4 265 57.46 42.19 35.62 32.58 16.42 36.85 Wanda 2:4 16.54 64.32 57.84 41.63 43.47 23.44 46.14 SparseGPT 2:4 13.48 68.58 56.72 45.36 46.89 25.92 48.69 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 Magnitude 2:4 969 62.16 48.74 43.48 41.57 20.63 43.32 Wanda 2:4 13.29 71.76 60.22 54.20 51.18 28.33 53.14 SparseGPT 2:4 11.55 74.21 60.77 57.25 53.03 29.44 54.94 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 Magnitude 2:4 1.2e4 64.38 52.19 49.46 44.58 21.25 46.37 Wanda 2:4 13.49 72.25 61.72 57.97 53.96 29.69 55.12 SparseGPT 2:4 11.17 74.16 62.43 59.18 56.27 31.74 56.76 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 Magnitude 2:4 54.59 64.42 60.93 56.72 59.18 27.31 53.71 Wanda 2:4 11.35 70.84 62.27 55.33 57.58 31.91 55.59 SparseGPT 2:4 8.67 72.14 64.96 58.93 60.90 34.22 58.23 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 Magnitude 2:4 8.33 70.47 62.04 50.65 62.46 31.74 55.47 Wanda 2:4 8.32 73.94 67.01 63.09 64.31 37.80 61.23 SparseGPT 2:4 7.07 75.46 68.51 65.52 66.04 39.76 63.06 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
-Appendix D Changes in the Parts Without Regularization
+Following previous works (Ashkboos et al.
+2024; Song et al.
+2024), for the comparison unstructured pruning methods like Mangnitude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by DReSS for regularization.
+Appendix E Optimal $\lambda$ for Different Models
+Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$.
+we list $\lambda$ for the best performance of each model in the Table 7:
+Model Phi-2 OPT-2.7B OPT-6.7B OPT-13B LLaMA2-7B LLaMA2-13B
+$\lambda$ $10^{-3}$ $5\times 10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$
+Table 7: The optimal $\lambda$ for each model.
+Appendix F Changes in the Parts Without Regularization
 As illustrated in Figure 6, the magnitude of the unregularized parameters exhibits an increase after the application of regularization, suggesting a redistribution of model capacity.
@@
 Figure 6: Ratio of the sum of absolute values of unregularized parameters after LLaMA2-7B regularization to the sum of absolute values of the corresponding parameters before regularization.
-For example, regularization is not applied to the first 75% of rows in $W_{Q}$, and the absolute sum of the rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Increased Ratios.
-This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the unregularized columns is compared to the original sum.
-Appendix E Performance of DReSS under Different Pruning Ratios and Datasets
-E.1 The Perplexity of DReSS under Different Pruning Ratios and Datasets
+For example, regularization is not applied to the first 75% of rows in $\mathbf{W_{Q}}$, and the absolute sum of the rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Increased Ratios.
+This is applied similarly to other matrices such as $\mathbf{W_{K},W_{V},W_{o},W_{gate},W_{up},W_{down}}$, and for layer activations, the sum of the unregularized columns is compared to the original sum.
+Appendix G Performance of DReSS under Different Pruning Ratios and Datasets
+G.1 The Perplexity of DReSS under Different Pruning Ratios and Datasets
 In Section 4.2, we utilize 1,000 samples randomly selected from the WikiText-2 dataset to guide the regularization process.
@@
 This highlights the scalability of DReSS and its potential to maintain model efficiency under varying levels of sparsity.
+Model Pruning Ratio WikiText-2 Alpaca PTB C4
+OPT-2.7B Dense 12.46 11.64 17.97 14.32
+10% 12.48 11.71 18.16 14.71
+20% 12.50 11.93 19.45 15.42
+30% 14.49 12.88 23.58 18.93
+40% 18.92 15.46 31.30 24.33
+50% 24.57 21.37 45.32 32.15
+60% 33.83 31.22 58.73 45.56
+OPT-6.7B Dense 10.85 10.27 15.77 12.71
+10% 10.91 10.45 16.05 13.18
+20% 11.02 10.83 17.58 14.45
+30% 12.32 11.46 19.65 16.68
+40% 14.26 12.71 25.52 20.94
+50% 19.63 15.66 33.78 28.22
+60% 28.75 21.39 47.29 39.47
+OPT-13B Dense 10.12 9.46 14.52 12.06
+10% 10.22 9.65 14.88 12.37
+20% 10.31 9.97 15.64 13.15
+30% 10.99 10.56 18.81 15.63
+40% 11.62 11.25 23.07 19.55
+50% 13.85 13.23 29.26 26.21
+60% 27.63 20.12 38.59 35.58
+LLaMA2-7B Dense 5.47 5.25 7.92 7.26
+10% 5.54 5.29 8.06 7.34
+20% 5.63 5.37 8.29 7.79
+30% 7.39 7.32 9.13 8.46
+40% 9.62 8.62 12.37 10.56
+50% 13.77 12.59 18.85 15.18
+60% 24.38 20.25 29.34 27.37
+LLaMA2-13B Dense 4.88 4.63 7.16 6.73
+10% 4.94 4.69 7.23 6.96
+20% 5.08 4.83 7.61 7.53
+30% 5.61 5.42 8.33 8.14
+40% 6.25 6.14 9.27 9.05
+50% 7.59 7.28 11.58 10.88
+60% 12.77 11.78 15.16 13.62
 Table 8: Perplexity comparison of DReSS with different pruning ratios.
@@
 For DReSS, we use the $\ell_{2}$-norm.
-Model Pruning Ratio WikiText-2 Alpaca PTB C4 OPT-2.7B Dense 12.46 11.64 17.97 14.32 10% 12.48 11.71 18.16 14.71 20% 12.51 11.93 19.45 15.42 30% 14.49 12.88 23.58 18.93 40% 18.92 15.46 31.30 24.33 50% 24.57 21.37 45.32 32.15 60% 33.83 31.22 58.73 45.56 OPT-6.7B Dense 10.85 10.27 15.77 12.71 10% 10.91 10.45 16.05 13.18 20% 11.02 10.83 17.58 14.45 30% 12.32 11.46 19.65 16.68 40% 14.26 12.71 25.52 20.94 50% 19.63 15.66 33.78 28.22 60% 28.75 21.39 47.29 39.47 OPT-13B Dense 10.12 9.46 14.52 12.06 10% 10.26 9.65 14.88 12.37 20% 10.47 9.97 15.64 13.15 30% 10.99 10.56 18.81 15.63 40% 11.62 11.25 23.07 19.55 50% 16.15 14.68 29.26 26.21 60% 27.63 20.12 38.59 35.58 LLaMA2-7B Dense 5.47 5.25 7.92 7.26 10% 5.54 5.29 8.06 7.34 20% 5.63 5.37 8.29 7.79 30% 7.39 7.32 9.13 8.46 40% 9.62 8.62 12.37 10.56 50% 13.77 12.59 18.85 15.18 60% 24.38 20.25 29.34 27.37 LLaMA2-13B Dense 4.88 4.63 7.16 6.73 10% 4.94 4.69 7.23 6.96 20% 5.08 4.83 7.61 7.53 30% 5.61 5.42 8.33 8.14 40% 6.25 6.14 9.27 9.05 50% 7.59 7.28 11.58 10.88 60% 12.77 11.78 15.16 13.62
-E.2 The Accuracy of DReSS under Different Pruning Ratios on Zero-shot Tasks
+G.2 The Accuracy of DReSS under Different Pruning Ratios on Zero-shot Tasks
 To systematically evaluate the performance of DReSS on zero-shot tasks under varying pruning rates, we adopt the experimental setup outlined in Section 4.2, where 1,000 samples are randomly selected from the WikiText-2 dataset to guide the regularization process.
@@
 Notably, the analysis reveals that DReSS maintains competitive performance even at higher pruning rates, demonstrating its effectiveness in preserving reasoning and commonsense understanding across different tasks.
+Model Pruning Ratio PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc
+OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39
+10% 70.38 60.12 51.79 52.46 29.78 52.91
+20% 69.96 59.47 50.45 51.87 28.62 52.07
+30% 64.52 58.76 48.25 50.23 27.52 49.86
+40% 61.32 56.27 47.39 48.26 25.57 47.76
+50% 59.38 53.46 44.63 46.19 21.66 45.06
+60% 52.99 48.74 40.03 43.35 16.25 40.27
+OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70
+10% 75.89 64.61 64.69 58.53 33.47 59.44
+20% 75.12 64.23 62.56 57.34 32.95 58.44
+30% 72.52 62.63 58.27 54.48 29.99 55.58
+40% 67.37 58.59 52.12 49.38 26.87 50.87
+50% 61.48 55.62 46.46 47.68 22.97 46.84
+60% 55.73 51.52 43.38 45.85 18.62 43.02
+OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79
+10% 75.46 64.69 68.56 61.79 35.58 61.22
+20% 74.78 64.61 67.25 61.66 35.43 60.75
+30% 72.62 63.26 65.69 59.64 32.57 58.76
+40% 68.67 61.49 62.74 55.98 29.26 55.63
+50% 62.19 56.46 58.55 52.15 24.53 50.78
+60% 57.43 52.72 51.23 47.62 21.05 46.01
+LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00
+10% 77.38 68.16 71.28 69.26 43.73 65.96
+20% 76.42 67.35 68.26 66.73 41.68 64.09
+30% 72.29 64.87 58.53 63.48 39.27 59.69
+40% 68.66 61.49 55.42 58.61 36.52 56.14
+50% 61.32 55.68 51.79 52.35 31.55 50.54
+60% 56.45 51.79 48.96 48.98 27.26 46.69
+LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76
+10% 79.35 72.15 77.42 76.92 48.57 70.88
+20% 78.27 71.98 76.89 75.43 47.62 70.04
+30% 75.49 69.73 73.54 72.87 45.41 67.41
+40% 73.56 64.46 67.75 68.45 41.28 63.10
+50% 68.73 59.82 62.48 60.12 36.14 57.46
+60% 62.25 56.27 59.93 53.37 29.77 52.32
 Table 9: Accuracy comparison of DReSS with different pruning ratios.
@@
 ‘Avg_Acc’ represents the average accuracy.
-Model Pruning Ratio PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39 10% 70.38 60.12 51.79 52.46 29.78 52.91 20% 69.96 59.47 50.45 51.87 28.62 52.07 30% 64.52 58.76 49.25 50.23 27.52 50.06 40% 61.32 56.27 47.39 48.26 25.57 47.76 50% 59.38 53.46 44.63 46.19 21.66 45.06 60% 52.99 48.74 40.03 43.35 16.25 40.27 OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70 10% 75.89 64.61 64.69 58.53 33.47 59.44 20% 75.12 64.23 62.56 57.34 32.95 58.44 30% 72.52 62.63 58.27 54.48 29.99 55.58 40% 67.37 58.59 52.12 49.38 26.87 50.87 50% 61.48 55.62 46.46 47.68 22.97 46.84 60% 55.73 51.52 43.38 45.85 18.62 43.02 OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79 10% 75.46 64.69 68.56 61.79 35.58 61.22 20% 74.78 64.61 67.25 61.66 35.43 60.75 30% 72.62 63.26 65.69 59.64 32.57 58.76 40% 68.67 61.49 62.74 55.98 29.26 55.63 50% 62.19 56.46 58.55 52.15 24.53 50.78 60% 57.43 52.72 51.23 47.62 21.05 46.01 LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00 10% 77.38 68.16 71.28 69.26 43.73 65.96 20% 76.42 67.35 68.26 66.73 41.68 64.09 30% 72.29 64.87 58.53 63.48 39.27 59.69 40% 68.66 61.49 55.42 58.61 36.52 56.14 50% 61.32 55.68 51.79 52.35 31.55 50.54 60% 56.45 51.79 48.96 48.98 27.26 46.69 LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76 10% 79.35 72.15 76.42 76.92 48.57 70.68 20% 78.27 71.98 73.20 75.43 47.62 69.30 30% 75.49 69.73 71.54 72.87 45.41 67.00 40% 73.56 64.46 67.75 68.45 41.28 63.10 50% 68.73 59.82 62.48 60.12 36.14 57.46 60% 62.25 56.27 59.93 53.37 29.77 52.32
-Appendix F More Details of the Dependency on Calibration Dataset
-When using Alpaca, WikiText-2, C4, and PTB as calibration and RFT data, the perplexity of various methods on Alpaca, C4, and PTB is shown as follows:
+Appendix H Dependency on Calibration Dataset
+Since DReSS relies on data-driven regularization, we investigate its dataset dependency.
+We evaluated perplexity of four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4.
+For fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2.
+As shown in Figure 7, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness.
+When using Alpaca, WikiText-2, C4, and PTB as calibration and RFT data, the perplexity of various methods on WikiText-2, Alpaca, C4, and PTB is shown as follows:
+Figure 7: Comparison of perplexity on Wikitext-2 using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
 Figure 8: Comparison of perplexity on Alpaca using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
@@
 Figure 10: Comparison of perplexity on PTB using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
+Appendix I Future Work and Limitations
+The “regularize-then-prune-then-finetune” paradigm proposed by DReSS is generalizable.
+In this paper, we only explore pruning at the channel level (i.e., rows or columns of parameter matrices).
+In future work, we plan to extend the pruning units to entire transformer layers.
+Using transformer layer as pruning unit is expected to yield more significant acceleration.